### v1

# Local Causal Discovery for Structural Evidence of Direct Discrimination

Abstract
Fairness is a critical objective in policy design and algorithmic decision-making. Identifying the causal pathways of unfairness requires knowledge of the underlying structural causal model, which may be incomplete or unavailable. This limits the practicality of causal fairness analysis in complex or low-knowledge domains. To mitigate this practicality gap, we advocate for developing efficient causal discovery methods for fairness applications. To this end, we introduce local discovery for direct discrimination (LD3): a polynomial-time algorithm that recovers structural evidence of direct discrimination. LD3 performs a linear number of conditional independence tests with respect to variable set size. Moreover, we propose a graphical criterion for identifying the weighted controlled direct effect (CDE), a qualitative measure of direct discrimination. We prove that this criterion is satisfied by the knowledge returned by LD3, increasing the accessibility of the weighted CDE as a causal fairness measure. Taking liver transplant allocation as a case study, we highlight the potential impact of LD3 for modeling fairness in complex decision systems. Results on real-world data demonstrate more plausible causal relations than baselines, which took 197$\times$ to 5870$\times$ longer to execute.
1 Introduction
Fairness holds fundamental importance in policy design and algorithmic decision-making, especially in high-stakes domains such as healthcare, insurance, and policing [85, 19]. Various criteria have been proposed to measure unfairness with respect to protected or sensitive attributes, such as gender and ethnicity [93]. Legal doctrines generally differentiate between direct discrimination and indirect or spurious forms of unfairness, codifying the notion that the mechanism of unfairness matters [6, 13]. However, fairness criteria based solely on statistical associations cannot disentangle the direct, indirect, and spurious mechanisms of unfairness [39, 50, 69], limiting their informativeness and actionability from a policy intervention standpoint. Consequently, there is a growing emphasis on applying causal reasoning to questions of fairness [39], shifting the focus from associations to interventions and counterfactual outcomes [5].
Causal fairness analysis provides a theoretical framework for disentangling the mechanisms of unfairness using the language of structural causal models (SCMs) [69]. The decomposition of observed parity gaps into direct, indirect, and spurious effects is a centeral objective [67]. As a result, causal fairness analysis can be framed as a special case of mediation analysis [90]. When the complete ground truth SCM is known, the various pathways of unfairness can be read directly from structural equations or the underlying causal graph (Table A.1). With this structural knowledge in hand, points of potential intervention are made clear and causal measures of unfairness are identifiable.
Previous works on effect estimation in causal fairness [98, 67, 69, 68] and mediation analysis [59, 89, 82, 62] generally assume prior knowledge of the true causal graph. In practice, prior structural knowledge is often incomplete, absent, or contentious for complex domains, even among experts [65]. Even with prior knowledge, mediation analysis is challenging territory. Identifiability in causal mediation has been the subject of complicated discussion and debate among theoreticians [62], raising barriers to entry for applied researchers [89]. Consequently, causal fairness analysis is not easily applied to systems with complex causal models, posing a gap between theory and practice.
When the true SCM is unknown, the underlying structure can be learned from observational data through causal discovery [83, 26]. Global discovery recovers the full causal graph or its equivalence class (e.g., [84, 16, 80, 64, 102]), providing mechanistic insights in the natural and social sciences [73, 46]. However, global discovery can pose modeling challenges, including the curse of dimensionality [100], high sample complexity [84, 71, 42], and exponential time complexity in unconstrained search spaces [84, 16, 17, 18]. Reliability is also an open problem, as global methods can sometimes disagree with expert-defined DAGs in complex domains [79, 65]. In the fairness setting, comparisons across global discovery methods can yield conflicting fairness conclusions [7]. Additionally, given an outcome of interest, global discovery is weighed down by a sort of wastefulness: the relations among all observed variables are learned, whether or not they are causally relevant to the outcome [12].
$X$$Y$$\mathbf{C}$$\mathbf{M}$ Standard Fairness Model
$X$$Y$$\mathbf{C}$$\mathbf{M}$ Direct Mechanism
$X$$Y$$\mathbf{C}$$\mathbf{M}$ Indirect Mechanism
$X$$Y$$\mathbf{C}$$\mathbf{M}$ Spurious Mechanism
Figure 1: The Standard Fairness Model (SFM) is a local subgraph around a protected attribute $X$ and outcome $Y$ (Definition 2.2) [69]. Variables that are irrelevant to fairness analysis are abstracted away, leaving only the set of mediators ($\mathbf{M}$) and the set of confounders ($\mathbf{C}$). Causal fairness analysis reveals whether the dotted paths exist. Direct, indirect, and spurious pathways are in solid red.
Local causal discovery [25, 21] aims to mitigate these inefficiencies by only learning the substructures relevant for specific target variables or downstream tasks, such as causal effect estimation [77, 14, 31, 48, 78] or feature selection [95, 96]. As the Standard Fairness Model (SFM; Figure 1) is itself a local subgraph nucleated around an attribute and outcome of interest [69], local causal discovery offers a natural framework for causal fairness analysis. However, task-specific local discovery algorithms are definitionally one-size-does-not-fit-all. Thus, existing methods may not be optimal for fairness tasks, particularly with respect to time efficiency in the nonparametric setting.
Contributions
To our knowledge, this work presents the first local causal discovery method specifically tailored for causal fairness analysis. We explicitly address the practicality gap in applying causal fairness theory to complex real-world domains with unknown SCMs, while ensuring implementability and state-of-the-art computational efficiency.
  1. 1.
    We introduce local discovery for direct discrimination (LD3), [^1] Code is available on GitHub: https://anonymous.4open.science/r/LD3-4440 a time efficient local causal discovery algorithm that assesses the structural direct criterion (SDC) for direct discrimination detection [69]. LD3 performs a linear number of independence tests with respect to variable set size, with asymptotic guarantees on correctness under sufficient assumptions.
  2. 2.
    We define a graphical criterion that is sufficient for identifying the weighted controlled direct effect (CDE) [59, 62], a qualitative indicator of direct discrimination [98]. We prove that this criterion is satisfied by the structural knowledge returned by LD3, enhancing the accessibility of this causal fairness measure.
  3. 3.
    As a real-world case study of causal fairness analysis for complex domains, we use LD3 to recover evidence of direct sex-based discrimination in liver transplant allocation in the United States. LD3 recovered more plausible causal relations than local and global baselines [38, 31], which performed 42 to 1021$\times$ more tests and took 197 to 5870$\times$ longer to run.
Paper Structure
Section 2 defines essential concepts in causal fairness analysis and direct effect estimation. In Section 3, we introduce LD3 and describe our theoretical and methodological contributions. Section 4 contextualizes our contributions in the broader literature. Section 5 presents numerical validation, and Section 6 highlights a real-world case study on liver transplant allocation.
2 Preliminaries
Let univariate random variables be denoted by capital letters (e.g., $X$), with specific values in lowercase (e.g., $X=x$). Multivariate random variables or sets are denoted by boldface capital letters (e.g., $\mathbf{X}$), with vector values in bold lowercase (e.g., $\mathbf{X}=\mathbf{x}$). Graphs or function sets are denoted by calligraphic letters (e.g., $\mathcal{F}$). Let $pa(\cdot)$, $an(\cdot)$, $ch(\cdot)$, and $de(\cdot)$ denote the parent, ancestor, child, and descendant sets for a variable in causal graph $\mathcal{G}$, respectively.
2.1 Causal Fairness Analysis
Causal fairness analysis can be framed in the language of SCMs [69]. Though the theory of SCMs and the potential outcomes framework are logically equivalent [24, 33, 8], the graphical intuition supplied by SCMs is particularly suited for causal fairness analysis.
Definition 2.1 (Structural causal model (SCM), Pearl [59]).
A SCM is a 4-tuple $\langle\mathbf{V},\mathbf{U},\mathcal{F},P(u)\rangle$ where $\mathbf{U}=\{U_{i}\}_{i=1}^{n}$ denotes a set of exogenous variables determined by factors outside of the model, $\mathbf{V}=\{V_{i}\}_{i=1}^{n}$ denotes a set of observed endogenous variables that are determined by variables in $\mathbf{U}\cup\mathbf{V}$, =$\mathcal{F}=\{f_{i}\}_{i=1}^{n}$ denotes a set of structural functions such that $V_{i}=f_{i}(pa(V_{i}),U_{i})$, and $P(u)$ is a distribution over the exogenous variables $\mathbf{U}$.
Every SCM corresponds to a causal graphical model. To facilitate causal fairness analysis, the true causal graph for an SCM can be projected onto the SFM (Figure 1).
Definition 2.2 (Standard Fairness Model (SFM), Plečko & Bareinboim [69]).
Let $\mathcal{G}=(\mathbf{V},\mathbf{E})$ be a causal graph with vertices $\mathbf{V}$ and edges $\mathbf{E}$. Let $\mathcal{G}_{SFM}$ be the projection of $\mathcal{G}$ onto the SFM. This projection is obtained by selecting a protected attribute-outcome pair $\{X,Y\}\subset\mathbf{V}$ and identifying sets $\mathbf{M},\mathbf{C}\subseteq\mathbf{V}\setminus\{X,Y\}$ that satisfy a specific set of relations with respect to $\{X,Y\}$. Thus, $\mathcal{G}_{SFM}$ is the graph induced by $\{X,Y,\mathbf{M},\mathbf{C}\}\subseteq\mathbf{V}$ that encodes the following assumptions:
  • $X$ is a protected attribute and $Y$ is an outcome of interest.
  • $\mathbf{M}$ is a set of mediators with respect to $X$ and $Y$.
  • $\mathbf{C}$ is a set of confounders with respect to $X$ and $Y$, which may share a latent confounder with $X$ (per the bidirected edge in Figure 1).
Note that $\mathbf{C}$ and $\mathbf{M}$ can be of arbitrary dimensionality, including zero. Probing whether the red paths in Figure 1 exist and estimating the associated effect sizes are central tasks in causal fairness analysis.
Structural Fairness Criteria
Multiple criteria have been proposed for evaluating fairness from graphical structures [69]. For the purposes of this work, we are fundamentally interested in detecting graphical signatures of direct discrimination. We say that an SCM is fair or unfair with respect to direct discrimination when the following structural conditions hold.
Definition 2.3 (Structural direct criterion (SDC), Plečko & Bareinboim [69]).
An SCM is fair in terms of direct discrimination if and only if the protected attribute is not a parent of the outcome.
$SDC=\mathbf{1}(X\in pa(Y))=\begin{cases}1\quad\text{if $X$ is a parent of $Y$,}\\ 0\quad\text{if $X$ is not a parent of $Y$.}\end{cases}$ (1)
2.2 Controlled Direct Effect (CDE)
The CDE is a valid test for identifying direct discrimination, as the true value is non-zero if and only if there is a direct path from the protected attribute to the outcome [98]. Let $X=x$ and $X=x^{*}$ be the exposure value corresponding to treatment and no treatment (i.e., the reference value), respectively.
Definition 2.4 (Controlled direct effect (CDE), Vanderweele [89]).
The CDE measures the expected change in outcome as the exposure changes given covariates $\mathbf{S}$, where mediators $\mathbf{M}$ are uniformly fixed to a constant value $\mathbf{m}$ over the population (as simulated by conditioning on $\mathbf{M}$ [61]).
$CDE(\mathbf{m})$ $\coloneqq\mathbb{E}[Y_{x,\mathbf{m}}-Y_{x^{*},\mathbf{m}}]=\sum_{\mathbf{s}}\big{[}\mathbb{E}[Y|x,\mathbf{s},\mathbf{m}]-\mathbb{E}[Y|x^{*},\mathbf{s},\mathbf{m}]\big{]}P(\mathbf{s}).$ (2)
Definition 2.5 (Identifiability conditions of the CDE, Vanderweele [89]).
The CDE is identifiable from observational data when, given covariates $\mathbf{S}$, the following holds for all values $X=x,\mathbf{M}=\mathbf{m}$:
  1. 1.
    There is no latent confounding of the exposure and outcome given covariates $\mathbf{S}$, i.e., $Y_{x,\mathbf{m}}\perp\!\!\!\perp X|\mathbf{S}$.
  2. 2.
    There is no latent confounding of the mediators and outcome given $\{X,\mathbf{S}\}$, i.e., $Y_{x,\mathbf{m}}\perp\!\!\!\perp\mathbf{M}|X,\mathbf{S}$.
Unlike the natural direct effect, CDE identification does not forbid the existence of confounders for $\{\mathbf{M},Y\}$ that are descended from $X$ (Appendix C in Pearl [62]). There exist pathological examples where adjustment for such confounders induces bias, but unbiased CDE estimates can still be obtained for such adjustment sets when given the appropriate estimator [66, 62, 35] (see Appendix B).
When interaction between the mediator and exposure is present, the CDE does not display effect constancy: the CDE will be non-unique, taking on as many values as there are levels of $\mathbf{M}$. Effect constancy is guaranteed for the CDE in linear systems, but not in nonlinear systems [62]. For the nonparametric setting, we avoid assumptions about the absence of interaction yet still wish to obtain a unique CDE. Thus, we define the weighted CDE as the following expectation over $\mathbf{M}$.
Definition 2.6 (The weighted CDE over $\mathbf{M}$).
For all values of covariates $\mathbf{S}$ and mediators $\mathbf{M}$,
$WCDE$ $\coloneqq\sum_{\mathbf{m}}\sum_{\mathbf{s}}\big{[}\big{[}\mathbb{E}[Y|x,\mathbf{s},\mathbf{m}]-\mathbb{E}[Y|x^{*},\mathbf{s},\mathbf{m}]\big{]}P(\mathbf{s})\big{]}P(\mathbf{m}).$ (3)
As the weighted CDE is a weighted sum over elements that are well-defined under the identifiability conditions in Definition 2.5, the weighted CDE is identifiable under Definition 2.5 as well.
As a Fairness Metric
As shown in [98], if $X$ has no direct path to $Y$ then $CDE(\mathbf{m})=0$. Thus, $CDE(\mathbf{m})\neq 0$ qualitatively implies direct discrimination. As the weighted CDE is a sum, it can equal zero when all elements are zero or when non-zero elements sum to zero. However, it can only be non-zero if at least one $CDE(\mathbf{m})\neq 0$. Thus, we encourage caution in interpreting zero values.
Selecting Mediators and Confounders for Weighted CDE Estimation
As shown in Pearl [59], the subset of mediators that are parents of $Y$ is sufficient for CDE identification. Similarly, it is sufficient to control only for confounders that are adjacent to $X$ or $Y$, as all backdoor paths definitionally enter into $X$ and $Y$. In this work, we opt for identifying confounders that are parents of $Y$. Note that direct effects are identifiable when the exact causal relations among the selected confounders and mediators are not known [82], which is supportive of our local discovery approach.
Definition 2.7 (Valid adjustment for the weighted CDE).
A valid adjustment set for weighted CDE estimation blocks (1) all backdoor paths for $\{X,Y\}$, (2) all backdoor paths for $\{\mathbf{M},Y\}$, and (3) all mediator paths for $\{X,Y\}$. This blocks non-causal association through confounding paths and indirect causal association through mediator paths, isolating the direct causal effect.
[…]
Table 1: A true partitioning exists for any $\mathbf{Z}$ (from [48]).
2.3 Mapping Causal Partitions to the Standard Fairness Model
To facilitate efficient structure learning, we leverage the causal partition taxonomy defined in [48] (Table 1). Given an exposure-outcome pair $\{X,Y\}$, any arbitrary variable set $\mathbf{Z}$ can be uniquely partitioned into exactly eight disjoint subsets (which may be empty) that are defined by the types of causal paths that they share with $X$ and $Y$ [^2] See [48] for formal definitions of these partitions and their proxies. Note that the notion of proxy variables is not relevant in this setting, as we are only concerned with true confounders and mediators. Narrowing the focus of structure learning to the relationships that are causally relevant to the exposure and outcome [12], this partition taxonomy provides convenient building blocks for local discovery. Mapping this partition taxonomy to the SFM, $\mathbf{Z}_{1}$ is a superset of confounders $\mathbf{C}$ and $\mathbf{Z}_{3}$ is a superset of mediators $\mathbf{M}$.
$X$$Y$$\mathbf{Z}_{1}$$\mathbf{Z}_{3}$$\mathbf{Z}_{4}$
Figure 2: The local subgraph identified by LD3. Here, members of $\mathbf{Z}_{1}$, $\mathbf{Z}_{3}$, and $\mathbf{Z}_{4}$ are parents of $Y$. Arrows   signify adjacency and   signify paths with arbitrarily many intervening variables. The SDC indicates whether the dotted edge from $X$ to $Y$ exists.
3 Local Discovery for Direct Discrimination (LD3)
Algorithm 1 provides a constraint-based discovery procedure for assessing the SDC and learning a valid adjustment set for weighted CDE estimation from observational data. The number of independence tests performed scales linearly with the number of nodes. Proofs of correctness are given in Appendix C.
Sufficient Conditions for Structure Learning
We assume causal Markov, faithfulness, and acyclicity. We do not impose parametric assumptions on causal functions nor distributional forms. As for all constraint-based methods, the independence test selected may impose its own parametric assumptions. When these are not well-justified, we recommend nonparametric tests (e.g., [28, 29, 99, 71]).
Asymptotic guarantees on partition label correctness, SDC correctness, and valid adjustment for unbiased weighted CDE estimation hold under Assumptions (A1) and (A2).
  1. (A1)
    $Y$ has no descendants in the observed variable set. This assumption is relatively straightforward to justify in settings where $Y$ is a terminal variable with respect to temporal ordering (e.g., when the outcome is death, or a decision made at a known time point).
  2. (A2)
    All parents of $Y$ are observed. Latent variables that are not parents of $Y$ are permissible.
Remark 3.1 (Notes on (A2)).
This assumption is a consequence of the sufficient conditions for identifiability of the CDE (Definition 2.5) [59, 62]. We exploit (A2) in our approach to evaluating the SDC [69], as it ensures that $X$ is conditionally $d$-separated from $Y$ when there is no direct path from $X$ to $Y$. It also ensures that no colliders in $\mathbf{Z}_{2\notin de(Y)}$ are retained for adjustment. If some parents of $Y$ are unobserved, the SDC might return a value of 1 and the weighted CDE might be non-zero even if there is no direct discrimination present.
We prove in Appendix C that latent variables not adjacent to $Y$ have no impact on correctness.
Theorem 3.2.
Latent variables that are not parents of $Y$ do not affect Algorithm 1.
Algorithm 1 LD3: Learning structural evidence of direct discrimination from observational data.
0:  Exposure $X$, outcome $Y$, variable set $\mathbf{Z}$, independence test, significance level $\alpha$.
0:  Adjustment set $\mathbf{A}_{\text{{DE}}}\coloneqq\mathbf{Z}_{1\in pa(Y)}\cup\mathbf{Z}_{3\in pa(Y)}\cup\mathbf{Z}_{4\in pa(Y)}$ and SDC results.
0:  (A1) $Y$ has no observed descendants and (A2) all parents of $Y$ are observed.
1:  $\mathbf{Z}^{\prime}\leftarrow\mathbf{Z}$
2:  for $\forall\;Z\in\mathbf{Z}^{\prime}$ do
3:     if $Z\perp\!\!\!\perp X\land Z\perp\!\!\!\perp Y$ then $Z\in\mathbf{Z}_{8}$
[…]
6:  $\mathbf{Z}^{\prime}\leftarrow\mathbf{Z}^{\prime}\setminus\mathbf{Z}_{8}\cup\mathbf{Z}_{7}\cup\mathbf{Z}_{4}$
7:  for $\forall\;Z\in\mathbf{Z}^{\prime}$ do
[…]
9:  for $\forall\;Z_{4}\in\mathbf{Z}_{4}$ do
[…]
11:  $\mathbf{A}_{\text{{DE}}}\leftarrow\mathbf{Z}_{1\in pa(Y)}\cup\mathbf{Z}_{3\in pa(Y)}\cup\mathbf{Z}_{4\in pa(Y)}$
12:  if $X\perp\!\!\!\perp Y|\mathbf{Z}_{1\in pa(Y)}\cup\mathbf{Z}_{3\in pa(Y)}$ then $SDC\leftarrow 0$
13:  else $SDC\leftarrow 1$
14:  return  $\mathbf{A}_{\text{{DE}}},SDC$
Causal Partitions in this Setting
Given (A1), there are no descendants of $Y$ in $\mathcal{G}$. Thus, there are no $\mathbf{Z}_{6}$ nor $\mathbf{Z}_{2\in de(Y)}$. However, there may be $\mathbf{Z}_{2\notin de(Y)}$. Thus, $\mathcal{G}$ can contain the following seven (potentially empty) causal partitions with respect to $\{X,Y\}$: $\mathbf{Z}_{1}$, $\mathbf{Z}_{2\notin de(Y)}$, $\mathbf{Z}_{3}$, $\mathbf{Z}_{4}$, $\mathbf{Z}_{5}$, $\mathbf{Z}_{7}$, and $\mathbf{Z}_{8}$. LD3 returns the subgraph pictured in Figure 2: all $\mathbf{Z}_{1}$, $\mathbf{Z}_{3}$, and $\mathbf{Z}_{4}$ that are directly adjacent to $Y$.
3.1 A Graphical Criterion for Identifying the Weighted CDE
Definition 3.3 (Graphical criterion for identifying the weighted CDE).
We define the set $\mathbf{A}_{\text{{DE}}}$, which contains all parents of the outcome. Under the causal partition taxonomy defined in [48],
$\mathbf{A}_{\text{{DE}}}\coloneqq\mathbf{Z}_{1\in pa(Y)}\cup\mathbf{Z}_{3\in pa(Y)}\cup\mathbf{Z}_{4\in pa(Y)}.$ (4)
Theorem 3.4 ($\mathbf{A}_{\text{{DE}}}$ is a valid adjustment set for the weighted CDE).
The adjustment set $\mathbf{A}_{\text{{DE}}}$ satisfies the identification conditions for the weighted CDE (Definition 2.5) by blocking (1) all backdoor paths for $\{X,Y\}$, (2) all backdoor paths for $\{\mathbf{M},Y\}$, and (3) all frontdoor paths for $\{X,Y\}$.
Intuition.$\mathbf{A}_{\text{{DE}}}$ contains only and exactly all the parents of $Y$: confounders for $\{X,Y\}$ that are adjacent to $Y$ ($\mathbf{Z}_{1\in pa(Y)}$), mediators for $\{X,Y\}$ that are adjacent to $Y$ ($\mathbf{Z}_{3\in pa(Y)}$), and all parents of $Y$ that are marginally independent of $X$ ($\mathbf{Z}_{4\in pa(Y)}$). Thus, at least one member of every backdoor and frontdoor path is in $\mathbf{A}_{\text{{DE}}}$. This provides conditional $d$-separation of $X$ and $Y$ if and only if there is no edge $X\to Y$. Full proofs are provided in Appendix C.
Remark 3.5 (The role of $\mathbf{Z}_{4\in pa(Y)}$).
Note that $\mathbf{Z}_{3}$ and $Y$ can be confounded by $\mathbf{Z}_{1}$, $\mathbf{Z}_{3}$, or $\mathbf{Z}_{4}$. Including $\mathbf{Z}_{4\in pa(Y)}$ in $\mathbf{A}_{\text{{DE}}}$ helps guarantee identifiability of the weighted CDE without requiring exact knowledge of confounding for $\mathbf{Z}_{3}$ and $Y$. As $\mathbf{Z}_{4}$ is not post-treatment, it is not forbidden from adjustment under the backdoor criterion [58]. It has been shown that adjusting for $\mathbf{Z}_{4}$ can reduce estimate variance in propensity score models [11], and conventional intuition dictates that adjustment should minimally constrain the exposure and maximally constrain the outcome [72]. Thus, there may be statistical benefits to including $\mathbf{Z}_{4\in pa(Y)}$, though we do not formally explore this here.
4 Related Works
Few works in causal discovery have centered on fairness objectives [9, 7]. To our knowledge, this work presents the first local causal discovery method that is specifically tailored for causal fairness analysis. Below, we summarize two lines of literature closely connected to this work.
Estimating Direct Causal Effects
Multiple estimands have been proposed for direct effect estimation, including the CDE [59], the NDE [59], and the counterfactual direct effect (Ctf-DE) [98]. The CDE is an interventional quantity, while the NDE and Ctf-DE are counterfactual quantities [62, 98]. The NDE is popular in causal fairness analysis [69] as it is guaranteed to admit a valid decomposition of the total causal effect into direct and indirect effects in the nonparametric setting [59]. However, when direct discrimination is the quantity of interest, the CDE is a valid measure [98]. Previously, the CDE has garnered more use in policy evaluation than the NDE [89]. In this work, we favor the weighted CDE as a direct discrimination metric in complex domains because it requires fewer untestable assumptions over the data generating process [82] and less prior structural knowledge than the NDE [59, 89]. This permits unbiased estimation in a wider range of structures, and a lower burden of causal discovery. We provide an extended justification in Appendix B.
Graphical Criteria for Direct Effect Identification
Sufficient and necessary conditions for unbiased identification vary by estimand, but generally require covariate adjustment [91]. Formal graphical rules can allow the valid adjustment set to be read directly from known causal structures, even in the presence of latent confounding [63, 49, 72]. Graphical criteria for valid adjustment have been proposed for diverse causal quantities, most famously the backdoor criterion for total effect estimation [58]. Often, multiple valid criteria exist per estimand. These can vary in the amount of prior structural knowledge that they require, and criteria that require less knowledge are easier to implement in practice [92, 30]. Shpitser & Pearl [81] present a complete graphical criterion for conditional interventional distributions. Pearl [59] defines a graphical criterion for identifying the NDE from observational data (Corollary 1). Discussion, debate, and confusion around the sufficient and necessary conditions for NDE identification have inspired a flurry of papers [37, 36], as summarized in Pearl [62]. Shpitser & VanderWeele [82] present a complete graphical criterion for the NDE that reconciles some debate, showing where certain common criteria coincide.
Graphical criteria for direct effect estimation are often challenging for practitioners to implement [82]. To increase the practicality of covariate adjustment in complex and low-knowledge domains, we advocate for the development of graphical criteria that are (1) relatively intuitive and practical to implement in applied research, and (2) are satisfied by knowledge that can be reliably learned through data-driven methods. Thus, we apply the causal partition taxonomy of [48] to provide local graphical intuition for fast and practical adjustment set discovery.
5 Empirical Validation
Baselines
We compare LD3 against global and local discovery baselines, with two evaluation methods per algorithm (one with post-processing adapted for this setting and one without, as described in Appendix D): PC (PC and PC) [38], [^3] https://github.com/keiichishima/pcalg LDP (LDP and LDPpa) [48], and LDECC (LDECC and LDECC) [31]. PC is a popular global discovery method that has spawned many extensions. LDP is asymptotically guaranteed to return a valid adjustment set for total effect estimation under the backdoor criterion [48]. LDECC returns the parents and children of a target variable [31]. All baselines are constraint-based to facilitate direct comparisons on the number of independence tests performed and evaluation with an oracle. Throughout, $n$ denotes data sample size (abbreviated with $k=1000$). All experiments were performed on a local MacBook Pro (M2 Pro Chip; 16 GB memory).
Figure 3: Mean runtime for the Asia and Sachs benchmarks from bnlearn ($\chi^{2}$ tests; $\alpha=0.001$; 95% confidence intervals in shaded regions). For Asia (either $\to$ dysp), $|\mathbf{Z}|=6$ and $|\mathbf{A}_{\text{{DE}}}|=1$. For Sachs (erk $\to$ akt), $|\mathbf{Z}|=11$ and $|\mathbf{A}_{\text{{DE}}}|=1$. See Figure E.2 for total independence tests performed. Raw data are reported in Tables E.2 and E.4.
Adjustment Set Correctness
As correct assessment of the SDC and unbiased estimation of the weighted CDE require that all parents of the outcome are correctly identified, we evaluated the ability of LD3 to recover the true parent set in randomly generated DAGs. Using an oracle independence test to simulate access to infinite data, we empirically validated LD3 on the adjacency matrics of random directed acyclic Erdős-Rényi graphs [23]. [^4] https://r.igraph.org/ LD3 perfectly recovered parent sets for randomly selected exposure-outcome pairs in 90 unique graphs. Ten graphs were generated per DAG size, with node counts in $[5,10,20,30,40,50,100,200,500]$. Parent F1, recall, and precision were 100%. Runtimes and total independence tests as node and edge cardinality scale are shown in Figure E.1.
Next, we assesed DAGs from the bnlearn Bayesian Network Repository [76]. Results for the Asia (Lung Cancer) dataset [45] are reported in Tables E.1 and E.2. Results for the Sachs protein expression dataset [74] are reported in Tables E.3 and E.4. For each benchmark, we selected one exposure-outcome pair for which SDC = 1 and SDC = 0. For Asia, LD3 and LDPpa had the highest F1 scores for both exposure-outcome pairs, with 100% F1 when SDC = 1 starting at $n=2.5k$ (Table E.1). For SDC = 0, LD3 improved with increasing sample size, while other methods decreased or peaked. [^5] See Table E.1 for elaboration. Notably, neither PC nor PC identified a single causal parent of the outcome. For Sachs, only LDECC attained 100% F1 for both pairs across all sample sizes, though LD3 and LDPpa were closely behind and still provided valid adjustment sets for the weighted CDE in all settings (Table E.3). The primary benefits of LD3 were efficiency and lax assumptions: the number of independence tests performed by LDECC and PC is worst-case exponential with respect to variable set size for random structures, LDP is worst-case quadratic, and LDP only guarantees correctness under the sufficient condition that at least one member of $\mathbf{Z}_{4}$ and $\mathbf{Z}_{5}$ are observed. Efficiency pitfalls for baselines are emphasized in Figure E.1 and Tables E.2 and E.4.
Weighted CDE Estimation
To evaluate the ability of LD3 to accurately assess the SDC and enable weighted CDE estimation, we tested linear-Gaussian instantiations of an 18-node DAG with an M-structure (i.e., a member of $\mathbf{Z}_{2\notin de(Y)}$) [22], a mediator chain, and a member of $\mathbf{Z}_{4}$ that confounds a mediator and $Y$ (Figure D.1). The weighted CDE was estimated using linear regression. $\mathbf{A}_{\text{{DE}}}$ F1, precision, and recall and SDC accuracy converged toward 100% as sample size increased, with the weighted CDE converging toward the true direct effect with low variance (Figures E.3, E.4).
6 Real-World Case Study: Fairness in Liver Transplant Allocation
We apply our proposed algorithm to a real-world case study in healthcare and compare its findings with past studies to illustrate the value that LD3 can bring in practice.
Allocation Under Scarcity
Liver transplantation is a critical therapeutic option for patients with end-stage chronic liver disease and acute liver failure. In the US, demand for liver transplants significantly surpasses the supply of available donor livers [55]. Patients who require a liver transplant are placed on a national waiting list managed by the United Network for Organ Sharing (UNOS). [^6] https://unos.org/ Upon availability of a donor liver, the distribution policy sorts waitlisted patients by multiple prioritization criteria, such as medical urgency, compatibility, and location, while balancing competing objectives such as perceived policy fairness [43]. The US liver transplant system has seen several key policy changes designed to optimize distribution and improve patient outcomes [57] (see Appendix F). The model for end-stage liver disease (MELD) is one facet of this policy, used to determine priority based on illness severity [51]. Despite efforts to improve fair allocation [40], it is widely recognized that US organ allocation suffers from disparities on the basis of race, sex, and socioeconomic status [101].
While a thorough assessment of unfairness in liver transplantation is out of scope, we take sex-based disparities as a case study. Over time, women have been less likely than men to receive a transplant [52, 27, 75, 1]. Women are less represented on liver transplant waitlists primarily due to lower chronic liver disease rates [75, 56]. Height differences, MELD exception points, and underestimated MELD scores have been raised as potential explanations for sex-based disparities [1]. Prior investigations measure these disparities via statistical association [53], which cannot disentangle causal mechanisms [39]. To the best of our knowledge, there have been limited results on sex-based disparities using US national data collected after March 2018 [20]. To address the lack of causal analysis, we apply LD3 to recover structural evidence of direct sex-based discrimination in transplant allocation.
Detecting Discrimination
We utilize the National Standard Transplant Analysis and Research (STAR) dataset [70] containing patient-level information on waitlisted transplant candidates, recipients, and donors from the Organ Procurement & Transplantation Network (OPTN) up to October 2022. We filter to include adult patients awaiting or having undergone a whole liver transplant from deceased donors in 2017-2019 and 2020-2022, keeping features relevant for discrimination discovery (Appendix F). Time frames were selected as stable periods without changes in the UNOS allocation policy. Estimation was performed with double machine learning [15]. [^7] Double machine learning was implemented with econml (https://econml.azurewebsites.net) using random forest classifiers with a 70%/30% train-test split (https://scikit-learn.org). We assumed that no confounders for $\{\mathbf{M},Y\}$ that are descended from the exposure open backdoor paths when conditioned on [66, 62, 35]. Under all experimental settings and time frames, results from LD3 suggest that the effects of sex on receiving a liver are not fully explained by observed confounding and mediation (Table 2). Given the observed variables, causal evidence of direct discrimination is detected (SDC = 1) and corroborated by non-zero weighted CDE estimates with significant $p$-values. In all settings, $\mathbf{A}_{\text{{DE}}}$ contains key factors used in the liver distribution policy, including initial MELD, initial age, region, and active exception case [70]. Consistent results validate the reliability of LD3 in this complex domain. Our results are concordant with prior observations that differences in body size and medical condition contribute to sex-based disparities [54], as weight and diagnosis are in $\mathbf{A}_{\text{{DE}}}$ in all settings. We also highlight that predicted $\mathbf{A}_{\text{{DE}}}$ indicate potential discrimination with respect to ethnicity, education, and payment method. These socioeconomic factors may warrant further investigation for future policy interventions. These preliminary results are only qualitative indicators that further fairness analyses should take place.
Unos Policy (2017–2019) Unos Policy (2020–2022) LD3 ($\alpha=0.01$) LD3 ($\alpha=0.05$) LD3 ($\alpha=0.01$) LD3 ($\alpha=0.05$) $\mathbf{A}_{\text{{DE}}}$ active exception case, diagnosis, education, ethnicity, exception type, initial MELD, initial age, region, weight active exception case, diagnosis, education, ethnicity, exception type, initial MELD, initial age, recipient blood type, region, weight active exception case, diagnosis, education, ethnicity, initial MELD, initial age, payment method, region, weight active exception case, diagnosis, education, ethnicity, initial MELD, initial age, payment method, region, weight SDC 1 1 1 1 WCDE 0.033 [0.016,0.049] 0.025 [0.008,0.042] 0.032 [0.018,0.047] 0.032 [0.018,0.047] $p$-value 0.000 0.003 0.000 0.000
Table 2: Results obtained with nonparametric $\chi^{2}$ independence tests and double machine learning. Total features is 18 (Appendix F). For time period 2017-2019, total observations is $n=21101$ patients. For time period 2020-2022, total observations is $n=22807$ patients. With no covariate control, SDC = 1 and the weighted CDE is 0.049 [0.033,0.066] ($p$-value = 0.000).
SDC and weighted CDE results for PC [38] and LDECC [31] are reported in Tables E.5 and E.6. PC and LDECC qualitatively agree with LD3 on both discrimination metrics, though at times with higher weighted CDE values. However, PC, LDECC, and LD3 had relatively low agreement in terms of $\mathbf{A}_{\text{{DE}}}$. Adjustment set cardinality was lower for PC and LDECC, and was zero for PC in 2017-2019 ($\alpha=0.05$) and one for LDECC the same period ($\alpha=0.01$). Most settings for PC and LDECC returned adjustment sets that omitted key expected variables, such as known policy criteria (e.g., initial MELD score, active exception case, and initial age). PC results for both significance levels show multiple untrustworthy directed edges (e.g., weight $\to$ sex, height $\to$ sex, blood type $\to$ age, education $\to$ height, body mass index $\to$ height), undermining the credibility of results. Likewise, LDECC predicts that the outcome has children, which is known to be untrue.
Efficiency Gains with LD3
Discovery with sequential independence tests is infamously slow. Relative to LD3, vanilla PC performed 479–1021$\times$ more tests and took 2295–5870$\times$ longer to execute on an Apple M2 Pro Chip (Table E.7). LDECC performed 42–984$\times$ more tests and took 197–5774$\times$ longer to run. Speedups of this magnitude will be impactful for researchers who require relatively slow nonparametric tests with large sample sizes (e.g., [28, 29, 99, 71]) and for high-dimensional feature spaces, especially when access to compute is limited. Speedups reported throughout this work meet or exceed those reported in limited work on hardware acceleration for embarrassingly parallel constraint-based discovery, comparing single-threaded CPUs, multi-threaded CPUs, and GPUs [97, 32, 10]. For nonparametric discovery to increase its real-world impact, it will need to efficiently handle increasingly large data within reasonable academic compute limits.
7 Conclusion
This work advocates for increased practicality in causal fairness pipelines, with a mind to computational efficiency and implementability. We propose a time-efficient and asymptotically correct local discovery method for identifying two qualitative measures of direct discrimination: the structural direct criterion and the weighted CDE. The number of independence tests performed is worst-case linear in the number of nodes, while comparable baselines are worst-case quadratic and exponential. We illustrate the utility of our approach on a real-world clinical fairness problem, providing causal structural evidence of direct sex-based discrimination in the national liver transplant allocation policy.
Limitations and Future Directions
Future work could extend LD3 to allow for $Y$ with descendants. The assumption that all parents of $Y$ are observed is sufficient but not necessary for CDE identification and could be replaced with a different criterion. The performance of LD3 will be constrained by the properties of the selected independence test, and any future improvements in statistical independence testing will be of great benefit to the causal discovery community. Future work could consider discovery for other fairness estimands, such as the natural effects [62] and path-specific effects [3].
Broader Impacts and Ethical Implications
Matters of unfairness and direct discrimination have real, profound impacts on human well-being and are often challenging to detect in real-world retrospective data. Researchers who deploy automated methods like LD3 should be prudent in interpreting and publicly communicating the statistical significance and uncertainty of their findings, whether evidence for or against discrimination is found.
Appendix
Appendix A Supplemental Figures and Tables
Disparate Treatment Disparate Impacts
Direct Causal Mechanism Indirect Causal Mechanism Spurious Association
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}X}$$Y$$Z$ ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}X}$$Y$$Z$ ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}X}$$Y$$Z$
$X\sim\text{Bern}(0.3)$ $X\sim\text{Bern}(0.3)$ $X\sim\text{Bern}(0.3+0.1Z)$
$Z\sim\text{Bern}(0.2+0.3{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}X})$ $Z\sim\text{Bern}(0.2+0.3{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}X})$ $Z\sim\text{Bern}(0.2)$
$Y\sim\text{Bern}(0.1+0.2{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}X})$ $Y\sim\text{Bern}(0.1+0.2Z)$ $Y\sim\text{Bern}(0.1+0.2Z)$
Table A.1: Example data generating processes under direct, indirect, and spurious structural mechanisms of unfairness. Adapted from Bareinboim & Plecko [4]. Let $X$ denote a protected attribute, $Y$ an outcome of interest, and $Z$ a third measured variable.
Appendix B Further Discussion on Direct Effect Estimation
B.1 Comparing the Weighted Controlled Direct Effect to the Natural Direct Effect
We provide additional discussion on the weighted CDE and its most popular counterpart, the NDE. Recall that the CDE and weighted CDE are defined as follows:
$CDE(\mathbf{m})$ $\coloneqq\mathbb{E}[Y_{x,\mathbf{m}}-Y_{x^{*},\mathbf{m}}]$ (5)
$=\mathbb{E}[Y|do(X=x,\mathbf{M}=\mathbf{m})]-\mathbb{E}[Y|do(X=x^{*},\mathbf{M}=\mathbf{m})]$ (6)
$=\sum_{\mathbf{s}}\big{[}\mathbb{E}[Y|x,\mathbf{s},\mathbf{m}]-\mathbb{E}[Y|x^{*},\mathbf{s},\mathbf{m}]\big{]}P(\mathbf{s}).$ (7)
$WCDE$ $\coloneqq\sum_{\mathbf{m}}\sum_{\mathbf{s}}\big{[}\big{[}\mathbb{E}[Y|x,\mathbf{s},\mathbf{m}]-\mathbb{E}[Y|x^{*},\mathbf{s},\mathbf{m}]\big{]}P(\mathbf{s})\big{]}P(\mathbf{m}).$ (8)
Definition B.1 (Average NDE, Pearl [62]).
The NDE measures the expected change in outcome as the exposure changes, with mediators fixed at the values that they would have been per individual prior to this change in exposure. The NDE captures the portion of the total causal effect that would be transmitted to the outcome if the mediators did not respond to changes in the exposure.
For values of the exposure $x$ and $x^{*}$, values $\mathbf{s}$ of covariates $\mathbf{S}$, and values $\mathbf{m}$ of mediators $\mathbf{M}$, the average NDE is
$NDE(x,x^{*};Y)$ $=\mathbb{E}[Y_{x,\mathbf{M}_{x}^{*}}-Y_{x^{*},\mathbf{M}_{x}^{*}}].$ (9)
The average NDE in Markovian models is identifiable from observational data per Corollary 2 in Pearl [59], allowing counterfactual expressions to be replaced by probabilistic estimands:
$NDE(x,x^{*};Y)$ $=\sum_{\mathbf{s}}\sum_{\mathbf{m}}\big{[}\big{[}\mathbb{E}[Y|x,\mathbf{m}]-\mathbb{E}[Y|x^{*},\mathbf{m}]\big{]}P(\mathbf{m}|x^{*},\mathbf{s})\big{]}P(\mathbf{s}).$ (10)
We note that the weighted CDE and NDE differ only by the way that we weigh the CDE: for the weighted CDE, we weigh by $P(\mathbf{m})$, and for the NDE, we weigh by $P(\mathbf{m}|x^{*},s)$. This stems from the fact that the CDE is a function of interventional distributions, while the NDE is a counterfactual quantity. The NDE requires probabilities of nested counterfactuals and a cross-world counterfactual assumption [82, 60, 62, 2]. Nevertheless, Pearl [59] showed that the NDE can be expressed with do-notation and reduced to a “do-free” expression by valid covariate adjustment, allowing for identification and estimation [60]. Unlike the NDE, the CDE does not admit a valid decomposition of the total causal effect into direct and indirect effects for the general nonparametric setting [59]. Conversely, the total effect is always equal to the difference of the NDE and the reverse transition of the natural indirect effect (NIE), even in the presence of nonlinearities and exposure-mediator interactions [62]. However, in the absence of exposure-mediator interaction, the CDE and NDE are equivalent and the CDE can be used for total effect decomposition [89]. The guaranteed decomposability of the total effect into the NDE and NIE is a primary reason that the natural effects are employed in causal fairness analysis [69]. However, when only evidence of direct discrimination is needed, it suffices to estimate the CDE [98]. In practice, the CDE has traditionally been in more frequent use in policy evaluation settings than the NDE [89].
The present work is centrally concerned with direct discrimination only. For our current objectives, we favor the weighted CDE as a metric because it requires fewer assumptions over the data generating process and requires less prior structural knowledge than the NDE [89]. For example, the NDE requires that no confounder for $\{\mathbf{M},Y\}$ is descended from $X$ [62], as the same covariates that deconfound $\{\mathbf{M},Y\}$ must also be used to identify the counterfactual quantity $P(\mathbf{m}|x^{*},s)$. As explained in Appendix C of Pearl [62], if this covariate set contains a variable that is descended from the exposure, $P(\mathbf{m}|x^{*},s)$ is not identifiable even under random treatment assignment. While adjusting for confounders for $\{\mathbf{M},Y\}$ that are descended from the exposure can induce bias in CDE estimation, this bias is addressable when given an appropriate estimator [66, 62, 35]. Therefore, the presence or absence of such variables is something that can be assumed at the estimation stage at the discretion of the user, but does not impact the behavior of LD3. The absence of confounders of this form is challenging to verify [89], and may be unrealistic in real-world data. This limits the range of structures for which NDE estimates can be obtained. Additionally, the identifiability of the NDE requires that $\mathbf{S}$ also deconfounds $\{X,\mathbf{M}\}$, adding an extra burden of graphical knowledge. See Andrews & Didelez [2] for a discussion of the cross-world counterfactual assumption in NDE identification. Though the NDE has useful properties for certain forms of analysis, we restrict our attention to the weighted CDE in the present work for these reasons.
Appendix C Proofs
C.1 Correctness of Algorithm 1
Proof.
We show that each set of tests is correct when all stated assumptions are met.
  • $\mathbf{Z}_{8}$ (Line 3): As proven in [48], the independence relations tested in Line 3 uniquely define partition $\mathbf{Z}_{8}$ and are therefore an asymptotically correct test for identifying this partition, and this partition only.
  • $\mathbf{Z}_{5,7}$ (Line 4): As proven in [48], the independence relations tested in Line 4 define partition $\mathbf{Z}_{7}$ in all structures and $\mathbf{Z}_{5}$ in structures for which $|\mathbf{Z}_{1}|=0$ (i.e., no confounders, latent nor observed, exist for $X$ and $Y$). Therefore, this test is asymptotically guaranteed to return all $\mathbf{Z}_{7}$ for arbitrary structures, and all $\mathbf{Z}_{5}$ if $\mathbf{Z}_{1}$ is the empty set.
  • $\mathbf{Z}_{4}$ (Line 5): As proven in [48], the independence relations tested in Line 5 uniquely define partition $\mathbf{Z}_{4}$ and are therefore an asymptotically correct test for identifying this partition, and this partition only.
  • $\mathbf{Z}_{1\in pa(Y)}\cup\mathbf{Z}_{3\in pa(Y)}$ (Line 8): At this stage, the only partitions remaining under consideration are $\mathbf{Z}_{1}$, $\mathbf{Z}_{2\notin de(Y)}$, $\mathbf{Z}_{3}$, and $\mathbf{Z}_{5}$ (assuming that $\mathbf{Z}_{1}$ is not the empty set). By definition, no $\mathbf{Z}_{5}$ will ever be adjacent to $Y$, as this partition cannot share a causal path to $Y$ that is not mediated by $X$ [34, 47, 48]. All $\mathbf{Z}_{2\notin de(Y)}$ are descended from $\mathbf{Z}_{1}$, $\mathbf{Z}_{3}$, and/or $\mathbf{Z}_{4}$ (i.e., any partition that can enter into $Y$), but not from $Y$ itself. Thus, all $\mathbf{Z}_{2}$ and $\mathbf{Z}_{5}$ will be eliminated by eliminating variables not adjacent to $Y$ given $X\cup\mathbf{Z}_{4}\cup\{\mathbf{Z}^{\prime}\setminus Z\}$. This test will also result in the elimination of $\mathbf{Z}_{1}$ and $\mathbf{Z}_{3}$ that are not adjacent to $Y$, including proxy confounders and proxy mediators as defined in Maasch et al. [48].
  • $\mathbf{Z}_{4\in pa(Y)}$ (Line 10): Any intervening node on the causal path from $\mathbf{Z}_{4}$ to $Y$ ($\mathbf{Z}_{4}\to\cdots\to Y$) must also belong to a partition that can enter into $Y$ ($\cdots\to Y$). By definition, this is strictly $\mathbf{Z}_{1}$, $\mathbf{Z}_{3}$, and $\mathbf{Z}_{4}$. Therefore, to obtain $d$-separation between $Y$ and any member of $\mathbf{Z}_{4}$ that is not adjacent to $Y$, a correct conditioning set will contain members of $\mathbf{Z}_{1}$, $\mathbf{Z}_{3}$ and $\mathbf{Z}_{4}$. It is therefore sufficient to condition all $\mathbf{Z}_{1}$ and $\mathbf{Z}_{3}$ that are already known to be adjacent to $Y$, unioned with all known $\mathbf{Z}_{4}$. Any variable that is marginally dependent on $Y$ given this conditioning set is a member of $\mathbf{Z}_{4}$ that is adjacent to $Y$.
  • $\mathbf{A}_{\text{{DE}}}$ (Line 11): By definition of $\mathbf{A}_{\text{{DE}}}$.
  • SDC (Line 12): By definition of the SDC (Definition 2.3), $SDC=\mathbf{1}(X\in pa(Y))$. If $X$ is not a parent of $Y$, it will always be $d$-separable from $Y$ when all their confounding and mediating paths are blocked. If $X$ is a parent of $Y$, $X$ and $Y$ will be always conditionally dependent given any conditioning set. Therefore, the SDC will evaluate to 0 if and only if $X\perp\!\!\!\perp Y|\mathbf{Z}_{1\in pa(Y)}\cup\mathbf{Z}_{3\in pa(Y)}$.
C.2 Proof of Theorem 3.2
We show that latent variables that are not parents of $Y$ do not impact the asymptotic guarantees of Algorithm 1.
Proof.
Maasch et al. [48] have shown previously that the marginal and conditional independence tests used to identify $\mathbf{Z}_{4}$ and $\mathbf{Z}_{8}$ are not affected by latent variables. Intuition for this follows from the fact that these tests only require knowledge of $\{X,Y\}$ and the candidate variable $Z$. In Algorithm 1, the test for $\mathbf{Z}_{5,7}$ also relies only on knowledge of $X$, $Y$, and $Z$, and is therefore correct in the presence of latent variables. As we do not require identification of $\mathbf{Z}_{5}$ itself for downstream results, this test suffices. Otherwise, latent variables could pose identifiability problems for $\mathbf{Z}_{5}$, as in prior works [48]. (A2) alone is sufficient to ensure $d$-separation between $Y$ and its non-parents at Lines 8 and 10, guaranteeing identification of the true parents of $Y$ in $\mathbf{Z}_{1}$, $\mathbf{Z}_{3}$, and $\mathbf{Z}_{4}$. Therefore, latent variables that are not parents of $Y$ have no impact on Algorithm 1. ∎
C.3 Proof of Theorem 3.4
Proof.
This theorem claims that $\mathbf{A}_{\text{{DE}}}\coloneqq\mathbf{Z}_{1\in pa(Y)}\cup\mathbf{Z}_{3\in pa(Y)}\cup\mathbf{Z}_{4\in pa(Y)}$ is a valid adjustment set for weighted CDE identificaiton. Per the identifiability conditions of the weighted CDE (Definition 2.5), a valid adjustment set must block all backdoor paths for $\{X,Y\}$ and $\{\mathbf{M},Y\}$. As shown in Pearl [59], knowledge of all mediators for $\{X,Y\}$ that are parents of $Y$ is sufficient for identifying the CDE (i.e., knowledge of every mediator in the full causal graph is unnecessary). Thus, we show that adjusting for $\mathbf{A}_{\text{{DE}}}$ blocks all backdoor paths for $\{X,Y\}$ and $\{\mathbf{M},Y\}$, does not open non-causal paths by conditioning on colliders, and blocks all frontdoor paths for $\{X,Y\}$ by conditioning on all mediators that are parents of $Y$ (Definition 2.6). We use the language of causal partitions defined in [48]. We assume that all assumptions of this work are met (e.g., (A2): all parents of $Y$ are observed in $\mathbf{Z}$).
  1. 1.
    Backdoor paths for $\{X,Y\}$ are blocked. Any confounder for $\{X,Y\}$ will be in causal partition $\mathbf{Z}_{1}$, by definition of $\mathbf{Z}_{1}$ (Table 1) [48]. All backdoor paths enter $Y$ by definition (i.e., $\cdots\to Y$). As $\mathbf{A}_{\text{{DE}}}$ contains all confounders that are parents of $Y$, all backdoor paths for $\{X,Y\}$ will be blocked when conditioning on $\mathbf{A}_{\text{{DE}}}$.
  2. 2.
    Backdoor paths for $\{\mathbf{M},Y\}$ are blocked. Any confounder for $\mathbf{M}$ and $Y$ will be in $\mathbf{Z}_{1}\cup\mathbf{Z}_{3}\cup\mathbf{Z}_{4}$, as these are the only causal partitions with edges pointing into $Y$. Set $\mathbf{A}_{\text{{DE}}}$ by definition contains all members of $\mathbf{Z}_{1}$, $\mathbf{Z}_{3}$, and $\mathbf{Z}_{4}$ that are parents of $Y$. Given (A2), all backdoor paths between $\mathbf{M}$ and $Y$ are guaranteed to be blocked.
  3. 3.
    Frontdoor paths are blocked. Any mediator for $\{X,Y\}$ will be in causal partition $\mathbf{Z}_{3}$, by definition of $\mathbf{Z}_{3}$ (Table 1) [48]. All frontdoor paths enter $Y$ by definition (i.e., $\cdots\to Y$). As $\mathbf{A}_{\text{{DE}}}$ contains all mediators that are parents of $Y$, all frontdoor paths will be blocked.
  4. 4.
    Non-causal paths through colliders are blocked. Though $\mathbf{Z}$ can contain colliders in $\mathbf{Z}_{2\notin de(Y)}$, $\mathbf{A}_{\text{{DE}}}$ does not contain members of $\mathbf{Z}_{2}$. Thus, non-causal paths will not be opened by conditioning on $\mathbf{A}_{\text{{DE}}}$. We show in Section C.1 that Algorithm 1 is asymptotically guaranteed to remove $\mathbf{Z}_{2\notin de(Y)}$ from the final adjustment set.
Appendix D Empirical Evaluation Methods
[…]
Figure D.1: Eighteen-node DAG used to evaluate the ability of LD3 to recover the true $\mathbf{A}_{\text{{DE}}}$, assess the SDC, and estimate the weighted CDE.
Baseline Evaluation
Baselines were evaluated on the basis of F1, precision, and recall with respect to the true members of $\mathbf{A}_{\text{{DE}}}$. Among baselines, only LDECC is designed to return the parents of $Y$. Additionally, both PC and LDECC return results relative to a Markov equivalence class (MEC). Therefore, we evaluate baselines both with and without post-processing to adjust for the task of $\mathbf{A}_{\text{{DE}}}$ prediction. This allows for out-of-the-box evaluation and a more charitable evaluation that is tailored for the task. In contrast, LD3 only needs to be run out-of-the-box. The two evaluation methods per algorithm are as follows: PC (PC and PC) [38], [^8] https://github.com/keiichishima/pcalg LDP (LDP and LDPpa) [48], and LDECC (LDECC and LDECC) [31]. We define the evaluation for each below.
  1. 1.
    PC: Without post-processing, we retain the parents of $Y$ in all DAGs in the MEC.
  2. 2.
    PC: With post-processing, we retain the parents of $Y$ in any DAG in the MEC.
  3. 3.
    LDECC: Without post-processing, we retain the variables labeled as parents of the target.
  4. 4.
    LDECC: With post-processing, we retain the variables labeled as parents of the target unioned with the variables that are unoriented.
  5. 5.
    LDP: Without post-processing, we retain $\mathbf{Z}_{1}\cup\mathbf{Z}_{\textsc{POST}}\cup\mathbf{Z}_{4}$. As our experimental setting requires that there are no descendants of the outcome, the superset $\mathbf{Z}_{\textsc{POST}}\coloneqq\mathbf{Z}_{2\in de(Y)}\cup\mathbf{Z}_{3}\cup\mathbf{Z}_{6}$ defined in Maasch et al. [48] will now only contain $\mathbf{Z}_{3}$.
  6. 6.
    LDPpa: We post-process the LDP output to only retain the variables that are not conditionally independent of $Y$ given the remaining members of the set.
Appendix E Empirical Results
Figure E.1: Total runtime and total conditional independence tests performed with respect to total edges and total nodes, using an oracle on random directed acyclic Erdős-Rényi graphs [23].
Figure E.2: Mean runtime for the Asia [45] and Sachs [74] benchmarks from bnlearn (https://www.bnlearn.com/bnrepository/) [76] ($\chi^{2}$ tests; $\alpha=0.001$; 95% confidence intervals in shaded regions). For Asia (either $\to$ dysp), $|\mathbf{Z}|=6$ and $|\mathbf{A}_{\text{{DE}}}|=1$. For Sachs (erk $\to$ akt), $|\mathbf{Z}|=11$ and $|\mathbf{A}_{\text{{DE}}}|=1$. Raw data are reported in Tables E.2 and E.4.
Figure E.3: Mean weighted CDE estimates over 100 replicates per sample size for linear-Gaussian instantiations of Figure D.1 (95% confidence intervals in shaded regions). Ground truth weighted CDE is denoted by the dashed line. LD3 used parametric Fisher-z tests ($\alpha=0.01$), as the data generating process was known for this simulation. Estimates were obtained with linear regression (https://scikit-learn.org/). Note: This pipeline is for ease of explication in the illustrative setting only. For unknown data generating processes, we do not recommend this pipeline due to restrictive parametric assumptions.
Figure E.4: SDC accuracy, $\mathbf{A}_{\text{{DE}}}$ F1, precision, and recall across sample sizes when SDC = 0 and SDC = 1. Values were averaged over 100 replicates of linear-Gaussian instantiations of Figure D.1, with 95% confidence intervals in shaded regions. LD3 used parametric Fisher-z tests ($\alpha=0.01$), as the data generating process was known for this simulation.
Asia (Lung Cancer) SDC = 1: either $\to$ dysp SDC = 0: xray ↛$\not\to$ dysp F1 Prec Rec F1 Prec Rec Oracle LD3 1.00 1.00 1.00 1.00 1.00 1.00 PC 1.00 1.00 1.00 1.00 1.00 1.00 PC 1.00 1.00 1.00 1.00 1.00 1.00 LDECC 1.00 1.00 1.00 1.00 1.00 1.00 LDECC 1.00 1.00 1.00 1.00 1.00 1.00 LDPpa 1.00 1.00 1.00 1.00 1.00 1.00 LDP 0.50 0.33 1.00 0.67 0.50 1.00 $n=2.5k$ LD3 1.00 [1.00,1.00] 1.00 [1.00,1.00] 1.00 [1.00,1.00] 0.80 [0.69,0.91] 1.00 [1.00,1.00] 0.70 [0.54,0.86] PC 0.00 [0.00,0.00] 0.00 [0.00,0.00] 0.00 [0.00,0.00] 0.00 [0.00,0.00] 0.00 [0.00,0.00] 0.00 [0.00,0.00] PC 0.00 [0.00,0.00] 0.00 [0.00,0.00] 0.00 [0.00,0.00] 0.00 [0.00,0.00] 0.00 [0.00,0.00] 0.00 [0.00,0.00] LDECC 0.75 [0.61,0.89] 0.65 [0.46,0.84] 1.00 [1.00,1.00] 0.67 [0.49,0.85] 0.65 [0.46,0.84] 0.70 [0.54,0.86] LDECC 0.75 [0.61,0.89] 0.65 [0.46,0.84] 1.00 [1.00,1.00] 0.67 [0.49,0.85] 0.65 [0.46,0.84] 0.70 [0.54,0.86] LDPpa 1.00 [1.00,1.00] 1.00 [1.00,1.00] 1.00 [1.00,1.00] 0.83 [0.72,0.94] 1.00 [1.00,1.00] 0.75 [0.59,0.91] LDP 0.26 [0.24,0.28] 0.15 [0.14,0.16] 0.97 [0.90,1.00] 0.14 [0.13,0.14] 0.07 [0.07,0.08] 1.00 [1.00,1.00] $n=5k$ LD3 1.00 [1.00,1.00] 1.00 [1.00,1.00] 1.00 [1.00,1.00] 0.90 [0.80,1.00] 1.00 [1.00,1.00] 0.85 [0.70,1.00] PC 0.00 [0.00,0.00] 0.00 [0.00,0.00] 0.00 [0.00,0.00] 0.00 [0.00,0.00] 0.00 [0.00,0.00] 0.00 [0.00,0.00] PC 0.00 [0.00,0.00] 0.00 [0.00,0.00] 0.00 [0.00,0.00] 0.00 [0.00,0.00] 0.00 [0.00,0.00] 0.00 [0.00,0.00] LDECC 0.60 [0.47,0.73] 0.47 [0.29,0.64] 1.00 [1.00,1.00] 0.52 [0.36,0.68] 0.47 [0.29,0.64] 0.60 [0.47,0.73] LDECC 0.60 [0.47,0.73] 0.47 [0.29,0.64] 1.00 [1.00,1.00] 0.52 [0.36,0.68] 0.47 [0.29,0.64] 0.60 [0.47,0.73] LDPpa 1.00 [1.00,1.00] 1.00 [1.00,1.00] 1.00 [1.00,1.00] 0.93 [0.85,1.00] 1.00 [1.00,1.00] 0.90 [0.77,1.00] LDP 0.65 [0.62,0.68] 0.48 [0.45,0.52] 1.00 [1.00,1.00] 0.65 [0.62,0.67] 0.48 [0.45,0.51] 1.00 [1.00,1.00] $n=10k$ LD3 1.00 [1.00,1.00] 1.00 [1.00,1.00] 1.00 [1.00,1.00] 0.90 [0.80,1.00] 1.00 [1.00,1.00] 0.85 [0.70,1.00] PC 0.00 [0.00,0.00] 0.00 [0.00,0.00] 0.00 [0.00,0.00] 0.00 [0.00,0.00] 0.00 [0.00,0.00] 0.00 [0.00,0.00] PC 0.00 [0.00,0.00] 0.00 [0.00,0.00] 0.00 [0.00,0.00] 0.00 [0.00,0.00] 0.00 [0.00,0.00] 0.00 [0.00,0.00] LDECC 0.50 [0.50,0.50] 0.33 [0.33,0.33] 0.00 [0.00,0.00] 0.40 [0.40,0.40] 0.33 [0.33,0.33] 0.50 [0.50,0.50] LDECC 0.50 [0.50,0.50] 0.33 [0.33,0.33] 0.00 [0.00,0.00] 0.40 [0.40,0.40] 0.33 [0.33,0.33] 0.50 [0.50,0.50] LDPpa 1.00 [1.00,1.00] 1.00 [1.00,1.00] 1.00 [1.00,1.00] 0.90 [0.80,1.00] 1.00 [1.00,1.00] 0.85 [0.70,1.00] LDP 0.67 [0.67,0.67] 0.50 [0.50,0.50] 1.00 [1.00,1.00] 0.64 [0.61,0.67] 0.47 [0.44,0.5] 1.00 [1.00,1.00]
Table E.1: Empirical performance on the Asia (Lung Cancer) Bayesian network dataset [44] from bnlearn [76]. Binary data were evaluated with $\chi^{2}$ independence tests ($\alpha=0.001$; 95% confidence intervals in brackets). Unexpected decreases in baseline performance for increasing sample size might be due to the difficulty of this dataset for structure learning, as noted in the bnlearn documentation: “Standard learning algorithms are not able to recover the true structure of the network because of the presence of a node (E) with conditional probabilities equal to both 0 and 1" (https://www.bnlearn.com/documentation/man/asia.html).
Asia (Lung Cancer) SDC = 1: either $\to$ dysp SDC = 0: xray ↛$\not\to$ dysp Time Tests Time Tests Oracle LD3 0.002 29 0.002 31 PC 0.007 110 0.007 115 LDECC 0.003 25 0.004 25 LDP 0.003 43 0.003 50 $n=2.5k$ LD3 0.003 [0.002,0.003] 23.4 [22.56,24.24] 0.003 [0.003,0.003] 27.2 [26.56,27.84] PC 0.032 [0.03,0.034] 67.1 [63.31,70.89] 0.033 [0.031,0.035] 70.5 [66.79,74.21] LDECC 0.024 [0.016,0.033] 35.5 [23.78,47.22] 0.018 [0.013,0.023] 35.5 [23.78,47.22] LDP 0.003 [0.003,0.003] 27.6 [24.41,30.79] 0.003 [0.003,0.003] 31.0 [27.94,34.06] $n=5k$ LD3 0.004 [0.003,0.004] 22.9 [21.96,23.84] 0.004 [0.003,0.004] 26.6 [26.0,27.2] PC 0.056 [0.054,0.057] 73.6 [72.02,75.18] 0.058 [0.057,0.06] 77.1 [75.29,78.91] LDECC 0.049 [0.044,0.054] 63.5 [56.21,70.79] 0.047 [0.042,0.053] 63.5 [56.21,70.79] LDP 0.073 [0.056,0.091] 214.4 [175.9,252.9] 0.039 [0.033,0.044] 192.0 [167.74,216.26] $n=10k$ LD3 0.004 [0.004,0.004] 22.0 [22.0,22.0] 0.005 [0.005,0.005] 26.6 [26.0,27.2] PC 0.100 [0.097,0.103] 77.4 [75.9,78.9] 0.105 [0.102,0.109] 82.3 [80.38,84.22] LDECC 0.096 [0.083,0.109] 67.1 [65.34,68.86] 0.087 [0.084,0.09] 67.1 [65.34,68.86] LDP 0.073 [0.056,0.091] 214.4 [175.9,252.9] 0.039 [0.033,0.044] 192.0 [167.74,216.26]
Table E.2: Total number of independence tests performaed and runtimes in seconds (95% confidence intervals in brackets).
Sachs Protein Expression SDC = 1: ERK $\to$ AKT SDC = 0: JNK ↛$\not\to$ P38 F1 Prec Rec F1 Prec Rec Oracle LD3 1.00 1.00 1.00 1.00 1.00 1.00 PC 1.00 1.00 1.00 1.00 1.00 1.00 PC 0.00 0.00 0.00 0.00 0.00 0.00 LDECC 1.00 1.00 1.00 1.00 1.00 1.00 LDECC 0.00 0.00 0.00 0.00 0.00 0.00 LDPpa 1.00 1.00 1.00 1.00 1.00 1.00 LDP 0.29 0.16 1.00 0.50 0.33 1.00 $n=5k$ LD3 0.97 [0.90,1.00] 0.95 [0.85,1.00] 1.00 [1.00,1.00] 0.96 [0.91,1.00] 0.93 [0.85,1.00] 1.00 [1.00,1.00] PC 0.00 [0.00,0.00] 0.00 [0.00,0.00] 0.00 [0.00,0.00] 0.77 [0.51,1.00] 0.80 [0.54,1.00] 0.75 [0.49,1.00] PC 0.00 [0.00,0.00] 0.00 [0.00,0.00] 0.00 [0.00,0.00] 0.73 [0.48,0.99] 0.80 [0.54,1.00] 0.70 [0.44,0.96] LDECC 1.00 [1.00,1.00] 1.00 [1.00,1.00] 1.00 [1.00,1.00] 1.00 [1.00,1.00] 1.00 [1.00,1.00] 1.00 [1.00,1.00] LDECC 0.10 [-0.1,0.30] 0.10 [-0.1,0.30] 0.10 [-0.1,0.30] 0.10 [-0.1,0.30] 0.10 [-0.1,0.30] 0.10 [-0.1,0.30] LDPpa 0.93 [0.85,1.02] 0.90 [0.77,1.00] 1.00 [1.00,1.00] 0.96 [0.91,1.00] 0.93 [0.85,1.00] 1.00 [1.00,1.00] LDP 0.29 [0.29,0.29] 0.17 [0.17,0.17] 1.00 [1.00,1.00] 0.50 [0.50,0.50] 0.33 [0.33,0.33] 1.00 [1.00,1.00] $n=10k$ LD3 1.00 [1.00,1.00] 1.00 [1.00,1.00] 1.00 [1.00,1.00] 1.00 [1.00,1.00] 1.00 [1.00,1.00] 1.00 [1.00,1.00] PC 0.50 [0.17,0.83] 0.50 [0.17,0.83] 0.50 [0.17,0.83] 1.00 [1.00,1.00] 1.00 [1.00,1.00] 1.00 [1.00,1.00] PC 0.00 [0.00,0.00] 0.00 [0.00,0.00] 0.00 [0.00,0.00] 0.47 [0.16,0.78] 0.50 [0.17,0.83] 0.45 [0.14,0.76] LDECC 1.00 [1.00,1.00] 1.00 [1.00,1.00] 1.00 [1.00,1.00] 1.00 [1.00,1.00] 1.00 [1.00,1.00] 1.00 [1.00,1.00] LDECC 0.10 [-0.1,0.30] 0.10 [-0.1,0.30] 0.10 [-0.1,0.30] 0.30 [0.00,0.60] 0.30 [0.00,0.60] 0.30 [0.00,0.60] LDPpa 1.00 [1.00,1.00] 1.00 [1.00,1.00] 1.00 [1.00,1.00] 1.00 [1.00,1.00] 1.00 [1.00,1.00] 1.00 [1.00,1.00] LDP 0.29 [0.29,0.29] 0.17 [0.17,0.17] 1.00 [1.00,1.00] 0.50 [0.50,0.50] 0.33 [0.33,0.33] 1.00 [1.00,1.00] $n=20k$ LD3 0.97 [0.90,1.00] 0.95 [0.85,1.00] 1.00 [1.00,1.00] 1.00 [1.00,1.00] 1.00 [1.00,1.00] 1.00 [1.00,1.00] PC 0.90 [0.70,1.00] 0.90 [0.70,1.00] 0.90 [0.70,1.00] 1.00 [1.00,1.00] 1.00 [1.00,1.00] 1.00 [1.00,1.00] PC 0.00 [0.00,0.00] 0.00 [0.00,0.00] 0.00 [0.00,0.00] 0.10 [-0.1,0.30] 0.10 [-0.1,0.30] 0.10 [-0.1,0.30] LDECC 1.00 [1.00,1.00] 1.00 [1.00,1.00] 1.00 [1.00,1.00] 1.00 [1.00,1.00] 1.00 [1.00,1.00] 1.00 [1.00,1.00] LDECC 0.00 [0.00,0.00] 0.00 [0.00,0.00] 0.00 [0.00,0.00] 0.00 [0.00,0.00] 0.00 [0.00,0.00] 0.00 [0.00,0.00] LDPpa 0.97 [0.90,1.00] 0.95 [0.85,1.00] 1.00 [1.00,1.00] 1.00 [1.00,1.00] 1.00 [1.00,1.00] 1.00 [1.00,1.00] LDP 0.29 [0.29,0.29] 0.17 [0.17,0.17] 1.00 [1.00,1.00] 0.50 [0.50,0.50] 0.33 [0.33,0.33] 1.00 [1.00,1.00]
Table E.3: Performance on the Sachs protein dataset [74] from bnlearn [76]. Discrete data were evaluated with $\chi^{2}$ independence tests ($\alpha=0.001$; 95% confidence intervals in brackets).
Sachs Protein Expression SDC = 1: ERK $\to$ AKT SDC = 0: JNK ↛$\not\to$ P38 Time Tests Time Tests Oracle LD3 0.013 37 0.013 37 PC 0.077 740 0.079 740 LDECC 0.079 653 0.081 646 LDP 0.017 46 0.017 52 $n=5k$ LD3 0.005 [0.005,0.006] 37.0 [37.0,37.0] 0.005 [0.005,0.005] 37.0 [37.0,37.0] PC 0.738 [0.729,0.747] 715.4 [708.08,722.72] 0.757 [0.741,0.773] 715.4 [708.08,722.72] LDECC 0.699 [0.688,0.709] 640.6 [638.15,643.05] 0.657 [0.596,0.718] 604.7 [550.92,658.48] LDP 0.006 [0.006,0.006] 47.2 [45.63,48.77] 0.009 [0.004,0.015] 52.4 [51.88,52.92] $n=10k$ LD3 0.007 [0.007,0.007] 37.0 [37.0,37.0] 0.007 [0.006,0.007] 37.0 [37.0,37.0] PC 1.388 [1.374,1.402] 730.4 [726.86,733.94] 1.394 [1.384,1.405] 730.4 [726.86,733.94] LDECC 1.199 [1.045,1.353] 605.1 [528.14,682.06] 1.041 [0.83,1.253] 536.7 [428.01,645.39] LDP 0.008 [0.008,0.008] 46.0 [46.0,46.0] 0.008 [0.008,0.009] 52.0 [52.0,52.0] $n=20k$ LD3 0.011 [0.011,0.011] 37.0 [37.0,37.0] 0.011 [0.01,0.011] 37.0 [37.0,37.0] PC 2.718 [2.689,2.746] 733.6 [732.19,735.01] 2.72 [2.676,2.763] 733.6 [732.19,735.01] LDECC 2.5 [2.451,2.549] 647.4 [644.05,650.75] 2.472 [2.438,2.506] 644.4 [641.0,647.8] LDP 0.012 [0.012,0.013] 46.6 [45.42,47.78] 0.013 [0.013,0.013] 52.0 [52.0,52.0]
Table E.4: Total number of independence tests and runtimes (95% confidence intervals in brackets).
Unos Policy (2017–2019) Unos Policy (2020–2022) PC ($\alpha=0.01$) PC ($\alpha=0.05$) PC ($\alpha=0.01$) PC ($\alpha=0.05$) $\mathbf{A}_{\text{{DE}}}$ education, recipient blood type None active exception case, intial MELD, recipient blood type active exception case, initial age, recipient blood type, region SDC 1 1 1 1 WCDE 0.046 [0.03,0.063] 0.049 [0.033,0.066] 0.037 [0.023,0.050] 0.028 [0.015,0.041] $p$-value 0.000 0.000 0.000 0.000
Table E.5: SDC assessment and weighted CDE estimation with double machine learning using results from PC Algorithm [84], as implemented by Kalisch & Buhlmann [38]. We define $\mathbf{A}_{\text{{DE}}}$ using the union method (PC; Appendix D), though results for $\mathbf{A}_{\text{{DE}}}$ were exactly the same using PC.
Unos Policy (2017–2019) Unos Policy (2020–2022) LDECC ($\alpha=0.01$) LDECC ($\alpha=0.05$) LDECC ($\alpha=0.01$) LDECC ($\alpha=0.05$) $\mathbf{A}_{\text{{DE}}}$ payment method payment method, region active exception case, diagnosis, initial MELD, initial age, recipient blood type diagnosis, recipient blood type, region SDC 1 1 1 1 WCDE 0.049 [0.032,0.065] 0.048 [0.032,0.064] 0.032 [0.019,0.046] 0.032 [0.017,0.048] $p$-value 0.000 0.000 0.000 0.000
Table E.6: SDC assessment and weighted CDE estimation with double machine learning using results from LDECC using the union method for selection $\mathbf{A}_{\text{{DE}}}$ (LDECC; Appendix D) [31]. Results for $\mathbf{A}_{\text{{DE}}}$ were almost concordant using LDECC, with the exception of 2020-2022 when $\alpha=0.05$ (where the intersection method excluded diagnosis).
LD3 LDECC PC Ratio PC/LD3 Ratio LDECC/LD3 Policy $\alpha$ Time Tests Time Tests Time Tests Time Tests Time Tests 2017-2019 0.01 0.100 75 20.304 3303 237.324 35943 2373.24 479.24 203.04 44.04 0.05 0.101 75 19.947 3188 231.844 36512 2295.48 486.83 197.50 42.51 2020-2022 0.01 0.107 75 47.402 6959 465.278 63225 4348.39 843.00 443.01 92.79 0.05 0.090 71 519.707 69926 528.372 72517 5870.80 1021.37 5774.52 984.87
Table E.7: Total number of independence tests performed and runtimes in seconds for LD3 vs PC [38] and LDECC [31] on the liver transplant allocation dataset. While PC is worst-case exponential in time complexity [84], LDECC is worst-case exponential for some structures and polynomial for others [31]. This could explain the wide variation in LDECC runtimes across experimental settings.
Appendix F Liver Fairness
Additional Background
The US liver transplant system has seen several key policy changes designed to optimize organ distribution and improve patient outcomes. In 2002, the model for end-stage liver disease (MELD) was adopted to determine patients’ priority for liver transplants based on the severity of their illness [51]. In June 2013, UNOS implemented the Share 35 policy to require that the regional sharing of livers for patients with a MELD score of $\geq 35$ prioritizes the local sharing of livers to less severe patients with a MELD score of $<35$ [94]. In January 2016, the MELD-sodium (i.e. MELD-Na) score went into effect to improve the original MELD by incorporating serum sodium levels into the calculation [88]. In February 2020, the acuity-circle (AC) policy for distributing livers replaced the longstanding donation service area based policy [86]. The AC policy uses a series of concentric circles, centered on the donor hospital, to determine the distribution range for available livers. In July 2023, the calculation formula for MELD was updated to include additional variables including female sex and serum albumin [87]. The new model, known as MELD 3.0, was expected to further refine the sorting of transplant candidates based on their medical urgency for transplant [41].
Feature Selection
We use the following features in the case study. Except for the liver allocation outcome, all features are determined at time of registration or listing. Thus (A1) holds. These features include all the possible parents to the liver allocation outcome available in the OPTN dataset, thus satisfying (A2). For time period 2017-2019, total observations is $n=21,101$ patients. For time period 2020-2022, total observations is $n=22,807$ patients. Below, we provide the definition of each variable included in our causal fairness analysis. Summary statistics are reported in Table F.1.
  1. 1.
    Sex (exposure, protected attribute): Recipient sex.
  2. 2.
    Liver allocation (outcome): Did the candidate receive a liver transplant?
  3. 3.
    Recipient blood type: Recipient blood group at registration.
  4. 4.
    Initial age: Age in years at time of listing.
  5. 5.
    Ethnicity: Recipient ethnicity category.
  6. 6.
    Hispanic/Latino: Is the recipient Hispanic/Latino?
  7. 7.
    Education: Recipient highest educational level at registration.
  8. 8.
    Initial MELD: Initial waiting list MELD/PELD lab score.
  9. 9.
    Active exception case: Was this an active exception case?
  10. 10.
    Exception type: Type of exception relative to hepatocellular carcinoma (HCC).
  11. 11.
    Diagnosis: Primary diagnosis at time of listing.
  12. 12.
    Initial status: Initial waiting list status code.
  13. 13.
    Number of previous transplants: Number of prior transplants that the recipient received.
  14. 14.
    Weight: Recipient weight (kg) at registration.
  15. 15.
    Height: Recipient height at registration.
  16. 16.
    BMI: Recipient body mass index (BMI) at listing.
  17. 17.
    Payment method: Recipient primary projected payment type at registration.
  18. 18.
    Region: Waitlist UNOS/OPTN region where recipient was listed or transplanted.
Unos Policy (2017-2019) Female (n = 7679) Male (n = 13422) p-value Test Active exception case 0.36 (0.73) 0.48 (0.83) 7.241e-28 t-test Diagnosis 1 (PSC: Primary Sclerosing Cholangitis) 0.03 (0.18) 0.04 (0.2) 0.037 $\chi^{2}$ Diagnosis 6 (AHF: acute hepatic failure) 0.06 (0.23) 0.02 (0.15) 0.004 $\chi^{2}$ Diagnosis 7 (Cancer) 0.09 (0.28) 0.16 (0.37) 0.010 $\chi^{2}$ Height 161.9 (7.46) 175.91 (8.51) 0.000 t-test Initial MELD 20.5 (10.21) 18.83 (9.87) 1.588e-31 t-test Payment method 0.53 (0.5) 0.54 (0.5) 0.012 $\chi^{2}$ Recipient age 54.46 (12.42) 56.03 (10.74) 4.091e-22 t-test Weight 75.97 (18.53) 90.63 (19.59) 0.000 t-test Unos Policy (2020-2022) Female (n = 8574) Male (n = 14233) p-value Test Active exception case 0.39 (0.58) 0.43 (0.63) 3.629e-07 t-test Ethnicity 9 (Multiracial, non-hispanic) 0.01 (0.08) 0.0 (0.07) 0.022 Fisher’s exact Exception type 1 (Unknown) 0.29 (0.45) 0.28 (0.45) 0.022 $\chi^{2}$ Height 161.96 (7.8) 176.36 (8.25) 0.000 t-test Initial MELD 22.13 (10.52) 20.99 (10.48) 1.862e-15 t-test Initial status 0.04 (0.2) 0.02 (0.12) 2.858e-31 t-test Recipient age 53.83 (12.75) 54.74 (11.64) 3.059e-08 t-test Weight 75.76 (18.71) 91.1 (20.32) 0.000 t-test
Table F.1: Mean values (standard deviations) for features with statistically significant differences between males and females ($\alpha=0.05$). Summary statistics for all features are available on GitHub (https://anonymous.4open.science/r/LD3-4440).