### v3

# Local Causal Discovery for Structural Evidence of Direct Discrimination

Abstract
Identifying the causal pathways of unfairness is a critical objective for improving policy design and algorithmic decision-making. Prior work in causal fairness analysis often requires knowledge of the causal graph, hindering practical applications in complex or low-knowledge domains. Moreover, global discovery methods that learn causal structure from data can display unstable performance on finite samples, preventing robust fairness conclusions. To mitigate these challenges, we introduce local discovery for direct discrimination (LD3): a method that uncovers structural evidence of direct unfairness by identifying the causal parents of an outcome variable. LD3 performs a linear number of conditional independence tests relative to variable set size, and allows for latent confounding under the sufficient condition that all parents of the outcome are observed. We show that LD3 returns a valid adjustment set (VAS) under a new graphical criterion for the weighted controlled direct effect, a qualitative indicator of direct discrimination. LD3 limits unnecessary adjustment, providing interpretable VAS for assessing unfairness. We use LD3 to analyze causal fairness in two complex decision systems: criminal recidivism prediction and liver transplant allocation. LD3 was more time-efficient and returned more plausible results on real-world data than baselines, which took 46$\times$ to 5870$\times$ longer to execute.
1 Introduction
Fairness holds fundamental importance in policy design and algorithmic decision-making, especially in high-stakes domains such as healthcare and policing (Starke et al. 2022; Corbett-Davies et al. 2023). Various criteria have been proposed for measuring unfairness with respect to protected or sensitive attributes, such as gender and ethnicity (Verma and Rubin 2018). Legal doctrines generally differentiate between direct discrimination and indirect or spurious forms of unfairness, codifying the notion that the mechanism of unfairness matters (Barocas and Selbst 2016; Carey and Wu 2022). However, fairness criteria based solely on statistical associations cannot disentangle these mechanisms (Kilbertus et al. 2017; Makhlouf, Zhioua, and Palamidessi 2020), limiting their informativeness and actionability for policy interventions. Consequently, there is a growing emphasis on applying causal reasoning in fairness analysis (Kilbertus et al. 2017), shifting the focus from associations to interventions and counterfactual outcomes.
Causal fairness analysis (CFA) provides a theoretical framework for disentangling the mechanisms of unfairness using the language of structural causal models (SCMs). Previous works in CFA (Zhang and Bareinboim 2018; Plecko and Bareinboim 2023) and the allied field of mediation analysis (Pearl 2014; VanderWeele 2016) generally assume significant prior structural knowledge in order to decompose direct, indirect, and spurious effects. In practice, structural knowledge is often incomplete, absent, or contentious for complex domains, even among experts (Petersen et al. 2023). Furthermore, the identifiability of direct and indirect effects has been a topic of extensive debate among theoreticians (Pearl 2014), raising barriers to entry for applied researchers (Vanderweele 2011). Thus, existing methodologies in CFA can be challenging to apply in complex systems.
Among the many fairness measures proposed in CFA, the controlled direct effect (CDE) is a relatively straightforward qualitative indicator of direct discrimination that takes non-zero values only when the exposure is a direct cause of the outcome (Zhang and Bareinboim 2018). The CDE has often been favored in policy evaluation over alternative direct effect measures (Vanderweele 2011; VanderWeele 2013) as it is more interpretable for real-world interventions and requires fewer untestable assumptions and less prior structural knowledge (Pearl 2001; Shpitser and VanderWeele 2011). To increase the practicality of CFA in complex domains, we choose to focus on the CDE as a starting point for this work.
$X$$Y$$\mathbf{C}$$\mathbf{M}$Direct
$X$$Y$$\mathbf{C}$$\mathbf{M}$Indirect
$X$$Y$$\mathbf{C}$$\mathbf{M}$Spurious
Figure 1: The standard fairness model (SFM) is compactly represented as a local subgraph around protected attribute $X$ and outcome $Y$ (Plečko and Bareinboim 2024). Variables that are irrelevant to CFA are abstracted away, leaving confounders ($\mathbf{C}$) and mediators ($\mathbf{M}$). Directed edges represent the existence of active directed paths, and bidirected edges denote potential confounding. This work aims to identify direct mechanisms of unfairness in a data-driven way.
In low-knowledge domains, we can support CFA by learning causal structure directly from observational data. However, global causal discovery is challenging in finite data due to high sample complexity (Spirtes, Glymour, and Scheines 2001) and exponential time complexity in unconstrained search spaces (Chickering, Heckerman, and Meek 2004; Claassen, Mooij, and Heskes 2013). Learned causal graphs often disagree with expert knowledge in complex domains (Shen et al. 2020; Petersen et al. 2023), and can yield conflicting causal fairness conclusions in CFA (Binkytė et al. 2023 and Section 6 of this paper).
While global discovery learns the relations among all observed variables, local causal discovery only learns the substructures relevant for downstream tasks, such as causal effect estimation (Gupta, Childers, and Lipton 2023; Maasch et al. 2024; Shah, Shanmugam, and Kocaoglu 2024) or feature selection (Yu et al. 2021; Yu, Liu, and Li 2021). As the standard fairness model (SFM; Figure 1) is represented as a local subgraph (Plečko and Bareinboim 2024), local discovery offers a natural framework for CFA. However, since task-specific local discovery algorithms are definitionally one-size-does-not-fit-all, existing methods may not be optimal for fairness tasks.
Contributions
This work aims to increase the practicality of CFA for direct discrimination in complex domains with unknown causal graphs. Our contributions are three-fold.
  1. 1.
    Local discovery for direct discrimination (LD3). This local causal discovery method leverages the problem structure in CFA to efficiently detect graphical signatures of direct discrimination. [^1] Code on GitHub: https://github.com/jmaasch/LD3 LD3 discovers the parents of an outcome variable in a linear number of conditional independence tests with respect to variable set size.
  2. 2.
    A graphical criterion for the weighted controlled direct effect (WCDE). This criterion is sufficient to identify a valid adjustment set (VAS) for the WCDE, a qualitative indicator of direct discrimination. This criterion is satisfied by the knowledge returned by LD3.
  3. 3.
    Real-world fairness analysis. We deploy LD3 for two fairness problems: (1) racial discrimination in recidivism prediction and (2) sex-based discrimination in liver transplant allocation. LD3 recovered more plausible causal relations than local and global baselines, which performed 11–1021$\times$ more tests and took 46–5870$\times$ longer to run.
2 Preliminaries
Let capital letters denote univariate random variables (e.g., $X$), with their values in lowercase (e.g., $X=x$). Multivariate random variables or sets are denoted by boldface capital letters (e.g., $\mathbf{X}$), with vector values in bold lowercase (e.g., $\mathbf{X}=\mathbf{x}$). Graphs or function sets are denoted by calligraphic letters (e.g., $\mathcal{F}$). Let $pa(\cdot)$ and $de(\cdot)$ denote the parent and descendant sets for a variable in causal graph $\mathcal{G}$, respectively.
2.1 Causal Fairness Analysis
CFA can be framed in the language of SCMs and their graphical representations (Plečko and Bareinboim 2024).
Definition 1 (Structural causal model, Bareinboim et al. 2022).
An SCM is a 4-tuple $\langle\mathbf{V},\mathbf{U},\mathcal{F},p(\mathbf{u})\rangle$ where $\mathbf{U}=\{U_{i}\}_{i=1}^{n}$ denotes a set of exogenous variables determined by factors external to the model, $\mathbf{V}=\{V_{i}\}_{i=1}^{n}$ denotes a set of observed endogenous variables determined by $\mathbf{U}\cup\mathbf{V}$, =$\mathcal{F}=\{f_{i}\}_{i=1}^{n}$ denotes a set of structural functions such that $V_{i}=f_{i}(pa(V_{i}),U_{i})$, and $p(\mathbf{u})$ is the distribution over $\mathbf{U}$.
Every SCM corresponds to a causal graphical model. To facilitate CFA, the true causal graph for an SCM can be compactly represented using the SFM (Figure 1).
Definition 2 (Standard fairness model, Plečko and Bareinboim 2024).
Let $\mathcal{G}=(\mathbf{V},\mathbf{E})$ be a causal graph with vertices $\mathbf{V}$ and edges $\mathbf{E}$. Let $\mathcal{G}_{\mathrm{SFM}}$ be the projection of $\mathcal{G}$ onto the SFM, which is obtained by (1) selecting a protected attribute-outcome pair $\{X,Y\}\subset\mathbf{V}$ and (2) identifying sets $\mathbf{M},\mathbf{C}\subseteq\mathbf{V}\setminus\{X,Y\}$ that meet the following conditions:
  • $\mathbf{M}$ is the set of mediators with respect to $X$ and $Y$;
  • $\mathbf{C}$ is the set of confounders with respect to $X$ and $Y$.
Note that $\mathbf{C}$ or $\mathbf{M}$ can be the empty set, and confounding can exist among $\mathbf{C}$, $\mathbf{M}$, $X$, $Y$ (Figure 1).
Structural Fairness Criteria
Multiple criteria have been proposed for evaluating fairness from graphical structures. Here, we focus on a criterion for direct discrimination.
Definition 3 (Structural direct criterion (SDC), Plečko and Bareinboim 2024).
An SCM is fair with respect to direct discrimination if and only if the SDC evaluates to 0:
$SDC$ =$=\begin{cases}1\quad\text{if $X$ is a parent of $Y$,}\\ 0\quad\text{if $X$ is not a parent of $Y$.}\end{cases}$ (1)
2.2 Controlled Direct Effect
The CDE can be used to test for direct discrimination, as the true value is non-zero if and only if there is a direct path from the protected attribute to the outcome (Zhang and Bareinboim 2018). Let $X=x$ and $X=x^{*}$ be the exposure values corresponding to treatment and no treatment, respectively.
Definition 4 (CDE, Pearl 2014).
The CDE measures the expected change in outcome as the exposure changes when mediators $\mathbf{M}$ are uniformly fixed to a constant value $\mathbf{m}$:
$\mathrm{CDE}(\mathbf{m})$ $\coloneqq\mathbb{E}[Y|do(x,\mathbf{m})]-\mathbb{E}[Y|do(x^{*},\mathbf{m})].$ (2)
Definition 5 (Identifiability conditions of the CDE, Vanderweele 2011).
When we have access to a covariate set $\mathbf{S}$ that controls for confounding of both $\{X,Y\}$ and $\{\mathbf{M},Y\}$, the CDE is identifiable under the following conditions. Let $Y_{\check{x},\check{\mathbf{m}}}$ denote the value of $Y$ under $X=\check{x},\mathbf{M}=\check{\mathbf{m}}$.
  1. 1.
    There is no latent confounding of the exposure and outcome given $\mathbf{S}$, i.e., $Y_{\check{x},\check{\mathbf{m}}}\perp\!\!\!\perp X|\mathbf{S}$ for all $\check{x},\check{\mathbf{m}}$.
  2. 2.
    There is no latent confounding of the mediators and outcome given $\{X,\mathbf{S}\}$, i.e., $Y_{\check{x},\check{\mathbf{m}}}\perp\!\!\!\perp\mathbf{M}|X,\mathbf{S}$ for all $\check{x},\check{\mathbf{m}}$.
Then, we obtain $CDE(\mathbf{m})$ by [^2] When $\mathbf{S}$ is not sufficient for valid control of confounding for both $\{X,Y\}$ and $\{\mathbf{M},Y\}$, alternative formulae may be required (Vanderweele 2011; Pearl 2014).
$\sum_{\mathbf{s}}\big{(}\mathbb{E}[Y\mid x,\mathbf{s},\mathbf{m}]-\mathbb{E}[Y\mid x^{*},\mathbf{s},\mathbf{m}]\big{)}P(\mathbf{s}).$ (3)
Note that the subset of mediators that are parents of $Y$ is sufficient for CDE identification (Pearl 2001).
Comparison to Alternative Measures
Several estimands can capture direct effects, including CDE (Pearl 2001), natural direct effect (NDE; Pearl 2001), and counterfactual direct effect (Ctf-DE; Zhang and Bareinboim 2018). While CDE is an interventional quantity, NDE and Ctf-DE are counterfactual quantities. In this work, we favor CDE as it requires fewer untestable assumptions over the data generating process (Shpitser and VanderWeele 2011) and less structural knowledge than NDE and Ctf-DE (Pearl 2001; Vanderweele 2011; Zhang and Bareinboim 2018). The main objective of this work is to assess whether the protected attribute is a direct cause of the outcome. The CDE, NDE, and Ctf-DE are all non-zero if and only if $X$ directly points toward $Y$. Thus, we advocate for this simpler estimand for practicality. See Appendix A for an extended comparison of estimands.
Remark 1 (Assumptions on Confounding).
Unlike the NDE, CDE identification does not forbid the existence of confounders for $\{\mathbf{M},Y\}$ that are descended from $X$ (Pearl 2014). As done previously (Vanderweele and Vansteelandt 2009), this work assumes structures in which adjusting for confounders of $\{\mathbf{M},Y\}$ does not induce post-treatment bias, allowing us to identify the CDE with the expression provided in Equation 3. If this assumption does not hold, unbiased CDE estimates can still be obtained given alternative estimators (Petersen, Sinisi, and Van Der Laan 2006).
Weighted CDE
When interaction between the mediator and exposure exists, the value of CDE may vary across different mediator values (Pearl 2014). To avoid assumptions about interaction while still obtaining a unique estimate, we define the weighted CDE (WCDE) as the following expectation over $\mathbf{M}$. For simplicity, we assume discrete $\mathbf{M}$ [^3] Definition 6 generalizes to continuous $\mathbf{M}$ by integrating over its possible values. We note that LD3 operates independently of the variable type of $\mathbf{M}$.
Definition 6 (WCDE, Pearl 2000444Though unnamed, this expression is noted on page 131.).
We define WCDE as
$\sum_{\mathbf{m}^{\prime}}\big{(}\mathbb{E}[Y\mid do(x,\mathbf{m}^{\prime})]-\mathbb{E}[Y\mid do(x^{*},\mathbf{m}^{\prime})]\big{)}P(\mathbf{m}^{\prime}),$ (4)
where $\mathbf{M}^{\prime}\subseteq\mathbf{M}$ are parents of $Y$. Per Equation 3, WCDE is identifiable as
$\sum_{\mathbf{m}^{\prime}}\sum_{\mathbf{s}}\big{(}\mathbb{E}[Y|x,\mathbf{s},\mathbf{m}^{\prime}]-\mathbb{E}[Y|x^{*},\mathbf{s},\mathbf{m}^{\prime}]\big{)}P(\mathbf{s})P(\mathbf{m}^{\prime}).$ (5)
Definition 7 (VAS for the WCDE).
Given Definitions 5 and 6, a VAS for WCDE estimation blocks (1) all backdoor paths for $\{X,Y\}$, (2) all backdoor paths for $\{\mathbf{M},Y\}$, and (3) all mediator paths for $\{X,Y\}$.
As a Fairness Metric
Similar to CDE, WCDE indicates direct discrimination when its value is non-zero. However, a zero value does not guarantee the absence of direct discrimination, as different CDE values may cancel each other out. Thus, we encourage caution when interpreting zero values.
2.3 Mapping Causal Partitions to the SFM
The methods introduced in this work leverage the causal partition taxonomy defined in Maasch et al. (2024) (Table 1). Given an exposure-outcome pair $\{X,Y\}$, any arbitrary variable set $\mathbf{Z}$ can be uniquely partitioned into eight disjoint subsets (which may be empty) that are defined by the types of causal paths that they share with $X$ and $Y$. By focusing structure learning on relationships that are causally relevant to the exposure and outcome, this partition taxonomy provides practical building blocks for local discovery. Mapping this partition taxonomy to the SFM, $\mathbf{Z}_{1}$ are confounders $\mathbf{C}$ (and their proxies) and $\mathbf{Z}_{3}$ are mediators $\mathbf{M}$ (and their proxies). [^5] See Maasch et al. (2024) for more formal partition definitions. Proxy variables are not relevant in this setting, as these cannot be parents of $Y$ and are not returned by LD3. When referring to the union of multiple partitions, we use notation of the form $\mathbf{Z}_{1,3}\coloneqq\mathbf{Z}_{1}\cup\mathbf{Z}_{3}$.
[…]
Table 1: Adapted from Maasch et al. (2024).
3 Local Discovery for Direct Discrimination
In CFA, we can translate the fairness query “Is direct discrimination present?” into the graphical query “Is the protected attribute a parent of the outcome?” To answer this graphical query, we focus on two indicators of parentage: the SDC and WCDE, where the former can be directly answered by LD3 and the latter can be estimated by using the VAS returned by LD3.
The input to LD3 (Algorithm 1) is a variable set $\mathbf{Z}$ of unknown causal relation to the protected attribute $X$ and outcome $Y$. Instead of learning the causal graph, LD3 learns causal partition labels (Table 1). This local learning approach abstracts away structural information that is impertinent to direct discrimination detection, resulting in a computationally efficient discovery procedure. LD3 performs sequential CI tests to iteratively discover the partition label ($\widehat{\mathbf{Z}}$) of each variable in $\mathbf{Z}$. Leveraging these labels, LD3 uses a test of d-separation to evaluate SDC (Lines 12–13). By conditioning only on $\widehat{\mathbf{Z}}_{1,3\in pa(Y)}$, this CI test offers sample efficiency benefits relative to conditioning on all $\mathbf{Z}$. Additionally, LD3 returns a VAS for WCDE containing $pa(Y)\setminus X$.
Time Complexity
Constraint-based discovery methods are typically analyzed by the number of CI tests performed (Spirtes, Glymour, and Scheines 2001). For-loops at Lines 2–5, 7–8, and 9–10 of Algorithm 1 perform $O(|\mathbf{Z}|)$ tests each. All remaining lines perform a constant number of operations. Thus, the total number of CI tests is of $O(|\mathbf{Z}|)$, ensuring scalability in real-world CFA.
Algorithm 1 LD3
0:  Exposure $X$, outcome $Y$, variable set $\mathbf{Z}$, CI test of choice, significance level $\alpha$.
0:  Adjustment set $\mathbf{A}_{\text{{DE}}}$, SDC results.
0:  Sufficient conditions A1 and A2.
1:  $\mathbf{Z}^{\prime}\leftarrow\mathbf{Z}$
2:  for $\forall\;Z\in\mathbf{Z}^{\prime}$ do
3:     if $Z\perp\!\!\!\perp X\land Z\perp\!\!\!\perp Y$ then $Z\in\widehat{\mathbf{Z}}_{8}$
[…]
6:  $\mathbf{Z}^{\prime}\leftarrow\mathbf{Z}^{\prime}\setminus\widehat{\mathbf{Z}}_{8}\cup\widehat{\mathbf{Z}}_{5,7}\cup\widehat{\mathbf{Z}}_{4}$
7:  for $\forall\;Z\in\mathbf{Z}^{\prime}$ do
[…]
9:  for $\forall\;\widehat{Z}_{4}\in\widehat{\mathbf{Z}}_{4}$ do
[…]
11:  $\mathbf{A}_{\text{{DE}}}\leftarrow\widehat{\mathbf{Z}}_{1\in pa(Y)}\cup\widehat{\mathbf{Z}}_{3\in pa(Y)}\cup\widehat{\mathbf{Z}}_{4\in pa(Y)}$
12:  if $X\perp\!\!\!\perp Y|\widehat{\mathbf{Z}}_{1\in pa(Y)}\cup\widehat{\mathbf{Z}}_{3\in pa(Y)}$ then $SDC\leftarrow 0$
13:  else $SDC\leftarrow 1$
14:  return  $\mathbf{A}_{\text{{DE}}},SDC$
Sufficient Conditions for Structure Learning
We assume causal Markov, faithfulness, and acyclicity. We do not impose parametric assumptions on causal functions nor distributional forms. As for all constraint-based methods, the independence test selected may impose its own parametric assumptions. When these are not well-justified, we recommend nonparametric tests (e.g., Gretton et al. 2005, 2007).
Theorem 1.
Asymptotic guarantees on partitioning and SDC correctness hold under Assumptions A1 and A2. Given WCDE identifiability by Equation 5 (Remark 1), A1 and A2 are also sufficient for VAS discovery.
  1. A1
    $Y$ has no descendants in the observed variable set. This is satisfied when $Y$ is a terminal variable in the temporal ordering (e.g., when outcome is death, a policy or algorithmic decision made at a known time point, etc.).
  2. A2
    All parents of $Y$ are observed. Latent variables that are not parents of $Y$ are permissible. Thus, this is a milder condition than assuming causal sufficiency.
Proof of Theorem 1 is in Appendix B. Note that assumptions A1 and A2 are sufficient but not necessary. A1 has been previously used to facilitate parent and ancestor learning (Soleymani et al. 2022; Cai et al. 2023). LD3 learns causal partitions directly from data, without assumptions on temporal ordering other than A1 and $Y$ cannot cause $X$. Assumption A2 is a consequence of the sufficient conditions for CDE identifiability, as Definition 5 requires blocking all spurious and indirect paths into $Y$. A2 allows for unobserved variables that are not in $pa(Y)$, a more relaxed assumption than the causal sufficiency typically required in discovery (e.g., Spirtes, Glymour, and Scheines 2001; Zheng et al. 2018; etc.). We exploit A2 to evaluate the SDC, as it ensures that $X$ is conditionally $d$-separable from $Y$ when there is no direct path from $X$ to $Y$. It also ensures colliders ($\mathbf{Z}_{2\notin de(Y)}$) are removed from the adjustment set.
We empirically demonstrate failure modes and robustness to violations of A2 in Appendix D.1, showing that A2 is sufficient but not necessary. We prove in Appendix B that latent variables not adjacent to $Y$ do not impact correctness.
Theorem 2.
Latent variables that are not parents of $Y$ do not affect Algorithm 1.
$X$$Y$$\mathbf{Z}_{1}$$\mathbf{Z}_{3}$$\mathbf{Z}_{2}$$\mathbf{Z}_{4}$$\mathbf{Z}_{5}$$\mathbf{Z}_{7}$$\mathbf{Z}_{8}$?........(A)
$X$$Y$$\mathbf{Z}_{2}$$\mathbf{Z}_{1}$$\mathbf{Z}_{3}$$\mathbf{Z}_{4}$?(B)
Figure 2: LD3 assesses whether the edge $X\to Y$ exists. (A) Allowable partitions under A1 and A2. (B) Parents of $Y$ returned by LD3. Nodes are partition sets or subsets. Partition interrelations and latent confounding are abstracted away. Bidirected edge $Y\leftrightarrow\mathbf{Z}_{2}$ signifies $\mathbf{Z}_{2\notin de(Y)}$. Edges with $\cdots$ are paths of arbitrary length. Solid edges are adjacencies.
Figure 3: Baseline results for parent discovery in Sangiovese. Independence test count (Tests) is reported for constraint-based methods. Time is in seconds. Shaded regions denote 95% confidence intervals over ten replicates.
Causal Partitions in this Setting
Given A1, there are no descendants of $Y$ in $\mathcal{G}$ and therefore no $\mathbf{Z}_{6}$ nor $\mathbf{Z}_{2\in de(Y)}$. However, there may be $\mathbf{Z}_{2\notin de(Y)}$. Thus, $\mathcal{G}$ can contain the following seven (potentially empty) causal partitions with respect to $\{X,Y\}$: $\mathbf{Z}_{1}$, $\mathbf{Z}_{2\notin de(Y)}$, $\mathbf{Z}_{3}$, $\mathbf{Z}_{4}$, $\mathbf{Z}_{5}$, $\mathbf{Z}_{7}$, and $\mathbf{Z}_{8}$ (Figure 2.A). LD3 returns the partition subsets in Figure 2.B: all $\mathbf{Z}_{1}$, $\mathbf{Z}_{3}$, and $\mathbf{Z}_{4}$ that are directly adjacent to $Y$.
Remark 2 (The observed WCDE under violations of A2).
While A2 is sufficient but not necessary for WCDE identifiability, each backdoor and frontdoor path must be blocked by at least one observable variable. If no variables on such a path are measured, then no algorithm can identify the true WCDE. Users should note that the identifiability of any direct effect measure is fundamentally limited by the observability of backdoor and frontdoor paths.
3.1 A Graphical Criterion for the Weighted CDE
Under conditions where the WCDE is identifiable by Equation 5 (Remark 1), we propose the following criterion.
Definition 8 (Graphical criterion for identifying the WCDE).
Under the causal partition taxonomy defined in Maasch et al. (2024), we define the set $\mathbf{A}_{\text{{DE}}}$ that contains all parents of the outcome:
$\mathbf{A}_{\text{{DE}}}\coloneqq\mathbf{Z}_{1\in pa(Y)}\cup\mathbf{Z}_{3\in pa(Y)}\cup\mathbf{Z}_{4\in pa(Y)}.$ (6)
Theorem 3 ($\mathbf{A}_{\text{{DE}}}$ is a VAS for the WCDE).
$\mathbf{A}_{\text{{DE}}}$ is a valid adjustment set for the WCDE (Equation 5), satisfying the identification conditions in Definition 5.
Intuition.$\mathbf{A}_{\text{{DE}}}$ contains exactly all the parents of $Y$: confounders of $\{X,Y\}$ adjacent to $Y$ ($\mathbf{Z}_{1\in pa(Y)}$), mediators of $\{X,Y\}$ adjacent to $Y$ ($\mathbf{Z}_{3\in pa(Y)}$), and all parents of $Y$ that are marginally independent of $X$ ($\mathbf{Z}_{4\in pa(Y)}$). Thus, at least one member of every backdoor and frontdoor path is in $\mathbf{A}_{\text{{DE}}}$. This provides conditional $d$-separation of $X$ and $Y$ if and only if there is no edge $X\to Y$. Proof is in Appendix B.
Remark 3 (The role of $\mathbf{Z}_{4}$).
Note that $\mathbf{Z}_{3}$ and $Y$ can be confounded by $\mathbf{Z}_{1}$, $\mathbf{Z}_{3}$, or $\mathbf{Z}_{4}$ (Figure B.1). Including $\mathbf{Z}_{4\in pa(Y)}$ in $\mathbf{A}_{\text{{DE}}}$ helps guarantee the identifiability of WCDE without requiring exact knowledge of confounding for $\mathbf{Z}_{3}$ and $Y$.
Remark 4 (Variance and Actionability).
Definition 8 defines a VAS for the WCDE in the general setting, irrespective of A1 and A2. In settings where LD3 is used to obtain $\mathbf{A}_{\text{{DE}}}$, $\mathbf{Z}$ itself constitutes a VAS (per A1 and A2). However, adjusting for all $\mathbf{Z}$ is not advised for two primary reasons:
  1. 1.
    Adjusting for all $\mathbf{Z}$ risks unnecessary adjustment, which can inflate estimator variance under finite data (Schisterman, Cole, and Platt 2009). To support statistical efficiency, we follow intuition provided by prior theorems on VAS optimality with respect to asymptotic variance, which dictate exclusion of $\mathbf{Z}_{5}$, inclusion of $\mathbf{Z}_{4}$, and generally favor control for parents of $Y$ (Rotnitzky and Smucler 2020; Henckel, Perković, and Maathuis 2022).
  2. 2.
    Ignoring causal structure limits the informativeness and actionability of fairness conclusions. Forgoing structure learning is a missed opportunity to identify potential structural mechanisms that could be redressed through interventions (e.g., policy change).
A full discussion of this topic is in Appendix C. Sections 5 and 6 provide empirical support for points (1) and (2).
4 Related Works
Discovery for CFA
Few works in causal discovery have centered on fairness objectives (Binkytė et al. 2023). These include learning Suppes-Bayes causal networks with maximum likelihood estimation (Bonchi et al. 2017) and applying PC Algorithm (Zhang, Wu, and Wu 2017). To our knowledge, this work presents the first local causal discovery method that is specifically tailored for CFA.
Local Discovery of Direct Causes
Learning the direct causes of a target has primarily garnered interest in causal feature selection (Soleymani et al. 2022). Various Markov blanket (MB) learners have been proposed, though many cannot distinguish parents from children and/or spouses (Yu et al. 2020). MB-by-MB (Wang et al. 2014), Causal Markov Blanket (CMB; Gao and Ji 2015), and Local Discovery Using Eager Collider Checks (LDECC; Gupta, Childers, and Lipton 2023) are constraint-based methods that differentiate parents and children when unambiguous over the Markov equivalence class (MEC). Like the global algorithm PC (Spirtes, Glymour, and Scheines 2001), MB-by-MB, CMB, and LDECC have exponential time complexity with respect to variable set size in the worst case. Local Discovery by Partitioning (LDP; Maasch et al. 2024) causally partitions variables around $\{X,Y\}$ using a quadratic number of CI tests with respect to variable set size; results can be post-processed to identify parents of $Y$ if A1 is imposed.
5 Empirical Validation on Synthetic Data
Baselines
Baselines represent a range of approaches that are available as open-source Python implementations. Local baselines are MB-by-MB (Wang et al. 2014), LDECC (Gupta, Childers, and Lipton 2023), and LDP (Maasch et al. 2024). Global baselines are PC (Spirtes, Glymour, and Scheines 2001), DirectLiNGAM (Shimizu et al. 2011), and NOTEARS (Zheng et al. 2018). Extended baseline descriptions and post-processing procedures are given in Appendix D.2. Besides LDP, all baselines assume causal sufficiency. PC, MB-by-MB, and LDECC return results in terms of the MEC. DirectLiNGAM assumes an additive noise model. DirectLiNGAM and NOTEARS assume linearity. All experiments used an Apple MacBook (M2 Pro Chip; 12 CPU cores; 16G memory).
Figure 4: Predicted parent sets, SDC, and WCDE estimates for COMPAS. Exposure is race (R; red) and outcome is general recidivism risk decile score (DS; blue). Known parents of DS are in yellow. A = age; CD = charge degree; JF = juvenile felonies; JM = juvenile misdemeanors; PC = priors count; S = sex. WCDE is reported with $p$-value ($p$) and 95% confidence intervals in brackets. All methods used $\chi^{2}$ CI tests ($\alpha=0.05$). Full results are in Figures E.3E.5.
Parent Discovery
We evaluated whether LD3 can recover true parent sets using an oracle independence test on random exposure-outcome pairs in 90 unique directed acyclic Erdős-Rényi graphs (node counts in $[5...500]$) (Erdős, Rényi et al. 1960). [^6] https://r.igraph.org/ Parent F1, recall, and precision were 100%. Runtimes and total CI tests as node and edge cardinality scale are shown in Figure D.2. In Appendix D.4, we show for a linear-Gaussian SCM that WCDE estimates converged toward the true direct effect with low variance when adjusting for $\mathbf{A}_{\text{{DE}}}$ discovered with LD3 (Figures D.3, D.4).
All baselines were assessed on the Sangiovese benchmark from the bnlearn repository (Scutari 2010), a linear-Gaussian model of Tuscan grape production (Magrini, Di Blasi, and Stefanini 2017). Ten replicate datasets were sampled at $n=[250,500,1000]$. All constraint-based methods used Fisher-z tests ($\alpha=0.01$). DirectLiNGAM assumes non-Gaussian noise and was expected to underperform. LD3 was generally most performant across metrics, with LDP performing similarly (Figure 3, Table D.5). NOTEARS was significantly slower than other methods. Comparisons of LD3 to LDECC, LDP, and PC on Asia (Lauritzen and Spiegelhalter 1988) and Sachs (Sachs et al. 2005) benchmarks are in Tables D.6 and D.8. Runtime comparisons are in Figure D.2 and Tables D.7, D.9.
Estimator Variance in Finite Samples
As discussed in Remark 4, $\mathbf{Z}$ itself is a VAS under A1 and A2. However, adjusting for $\mathbf{Z}$ risks unnecessary adjustment, which can inflate the asymptotic variance of the causal effect estimator. We demonstrate the impacts of variance inflation as sample size scales in both a linear and nonlinear SCM (Figure D.5). Estimate variance using all $\mathbf{Z}$ was 7.8×7.8$\times$7.8 × to 9.6×9.6$\times$9.6 × higher than using $\mathbf{A}_{\text{{DE}}}$ in the linear SCM (Table D.2) and at least 12.6×12.6$\times$12.6 × higher in the nonlinear SCM (Table D.3).
VAS Interpretability
Structure learning can improve the interpretability of the VAS by removing irrelevant variables. In some structures, this removal can substantially reduce VAS size. On two bnlearn benchmarks with moderate to large DAGs and small $\mathbf{A}_{\text{{DE}}}$ cardinality, discovery with LD3 reduced adjustment set cardinality by at least 95.4% relative to retaining all $\mathbf{Z}$ (Table D.4).
6 Real-World Causal Fairness Analyses
We deploy LD3 for two causal fairness settings: (1) racial discrimination in criminal recidivism prediction and (2) sex-based discrimination in healthcare. We compare the results of LD3 to PC and LDECC on the basis of $\mathbf{A}_{\text{{DE}}}$ quality and computational efficiency. Causal discovery used $\chi^{2}$ CI tests and WCDE estimation used double machine learning (Chernozhukov et al. 2018). [^7] https://econml.azurewebsites.net Estimators used random forest classifiers with a 70% / 30% train-test split. [^8] https://scikit-learn.org We assumed A1 and A2. Data preprocessing is described in Appendix E. Note that all CFA results are only preliminary qualitative indicators, and further analyses should take place.
6.1 Race and COMPAS Recidivism Prediction
Background
We assessed the ability of LD3 to facilitate CFA on the COMPAS dataset from ProPublica.999github.com/propublica/compas-analysis Correctional Offender Management Profiling for Alternative Sanctions (COMPAS) is a commercial algorithm for case management and decision support used by the US criminal justice system to assess risk of recidivism (NorthPointe 2015). ProPublica’s landmark exposé on COMPAS found that African Americans are “almost twice as likely as whites to be labeled a higher risk but not actually re-offend” (Angwin et al. 2022).
We examined racial bias in the COMPAS General Recidivism Risk model. Due to data availability, we limited our analyses to the most represented racial groups (black and white). The algorithm’s developer states that the model directly considers prior criminal history and juvenile delinquency, among other factors (NorthPointe 2015, p. 27). Our data contained multiple indicators of criminal history and juvenile delinquency, so these were used to assess the quality of causal discovery in lieu of complete ground truth. We used three significance levels for independence testing ($\alpha=0.005,0.01,0.05$) to assess stability of results (Figures 4, E.3E.5). We selected 11 features with $n=6150$ observations (2454 white, 3696 black; see Appendix E). We explored two outcomes: (1) general recidivism decile score to probe bias in the COMPAS algorithm, and (2) actual two-year recidivism, to examine factors in real outcomes.
Figure 5: Predicted parent sets, SDC, and WCDE estimates for STAR liver data. Exposure is patient sex (S; red) and outcome is receiving a liver (L; blue). Known parents of $L$ are in yellow. AE = active exception case; BT = recipient blood type; DX = diagnosis; ED = education; ET = ethnicity; EX = exception type; IA = initial age; IM = initial MELD; PM = payment method; RE = region; WE = weight. WCDE is reported with $p$-value ($p$) and 95% confidence intervals in brackets. All methods used $\chi^{2}$ CI tests ($\alpha=0.01$). Additional results are in Tables E.2E.4.
Results
At all significance levels, results from LD3, PC, and LDECC qualitatively agree that the effects of race on decile score are not fully explained by observed [^10] Note that if A2 was violated, some of the observed effects might be due to unobserved variables. This is unverifiable using ProPublica’s data alone. See Remark 2. confounding and mediation (SDC = 1) and that the WCDE is signficant ($p=0.000)$. LD3 successfully predicted that juvenile delinquency is a parent of decile score at all significance levels, while PC and LDECC never did. Age, priors count, and charge degree were nearly always predicted to be parents across methods and significance levels. For two-year recidivism, LD3 stably predicted that the WCDE of race was not significant ($p=0.87,0.68,0.77$).
In general, results for PC and LDECC were less stable than LD3. Both methods wavered between strong significance ($p=0.000$) and no significance ($p>0.4$) for the WCDE of race on two-year recidivism. LDECC parent sets were inconsistent depending on whether they were taken as the intersection or union across graphs in the MEC, requiring additional interpretation by the user. On average, PC and LDECC took 46$\times$ longer to run and performed at least 11.7$\times$ more tests than LD3 across experiments (Figure E.2).
6.2 Sex and Liver Transplant Allocation
Background
We apply LD3 to a real-world case study in the US healthcare system: fairness in liver transplant allocation. Liver transplantation is a critical therapeutic option for patients with end-stage chronic liver disease and acute liver failure. Demand significantly surpasses supply for donor livers (SRTR 2020), and patients are placed on a national waiting list managed by the United Network for Organ Sharing (UNOS). The distribution policy sorts waitlisted patients by multiple criteria, such as medical urgency, compatibility, and location (Latt, Niazi, and Pyrsopoulos 2022). Several key policy changes have sought to optimize distribution and improve patient outcomes (Papalexopoulos et al. 2023; see Appendix E.2), including the model for end-stage liver disease (MELD; Malinchoc et al. 2000). Despite efforts to increase fairness (Kim et al. 2022), it is widely recognized that US organ allocation suffers from disparities (Zhang et al. 2018).
We explored potential sex-based discrimination in liver allocation. Sex-based disparities have been observed as statistical associations (Mathur et al. 2011; Gordon and Ladner 2012; Sarkar et al. 2015; Oloruntoba and Moylan 2015; Allen et al. 2018; Nephew and Serper 2021), but have not been explored through a causal lens. We use the National Standard Transplant Analysis and Research (STAR) dataset (OPTN 2024) for adult patients during 2017-2019 ($n=21,101$) and 2020-2022 ($n=22,807$). See Appendix E.2 for feature selection and summary statistics.
Results
Under all experimental settings and time frames, results from LD3 suggest that the effects of sex on receiving a liver were not fully explained by observed confounding and mediation (Figure 5, Table E.2). Causal evidence of direct discrimination was detected (SDC = 1) and corroborated by non-zero WCDE with significant $p$-values ($<0.005$). In all settings, $\mathbf{A}_{\text{{DE}}}$ contains key factors used in the liver distribution policy, including initial MELD, initial age, region, and active exception case (OPTN 2024). Our results are concordant with prior observations that differences in body size and medical condition contribute to sex-based disparities (Nephew et al. 2017), as weight and diagnosis are in $\mathbf{A}_{\text{{DE}}}$ in all settings. Predicted $\mathbf{A}_{\text{{DE}}}$ also indicate potential discrimination with respect to ethnicity, education, and payment method. These socioeconomic factors may warrant further investigation for future policy interventions.
PC and LDECC qualitatively agree with LD3 on both discrimination metrics (Figure 5, Tables E.3, E.4). However, PC, LDECC, and LD3 had relatively low agreement in terms of $\mathbf{A}_{\text{{DE}}}$. Adjustment set cardinality was lower (and at times zero) for PC and LDECC. Most settings for PC and LDECC returned $\mathbf{A}_{\text{{DE}}}$ that omitted key expected variables, such as known policy criteria (e.g., initial MELD score, active exception case, and initial age). At both significance levels, PC returned multiple untrustworthy edges (e.g., weight $\to$ sex, height $\to$ sex, blood type $\to$ age, education $\to$ height, body mass index $\to$ height), undermining the credibility of results. Likewise, LDECC predicts that the outcome has children, which is known to be untrue. PC performed 479–1021$\times$ more tests and took 2295–5870$\times$ longer to execute than LD3 (Table E.5). LDECC performed 42–984$\times$ more tests and took 197–5774$\times$ longer to run.
7 Conclusion
This work advocates for increased practicality in causal fairness pipelines. We propose a time-efficient and asymptotically correct local discovery method for identifying two qualitative indicators of direct discrimination: the SDC and WCDE. For two real-world causal fairness analyses, LD3 returned more stable and plausible predictions with signficantly better computational efficiency relative to baselines.
Limitations and Future Directions
Future work could extend LD3 to allow for $Y$ with descendants. The assumption that all parents of $Y$ are observed is sufficient but not necessary for CDE identification and could be replaced with a different criterion. LD3 cannot differentiate $\mathbf{Z}_{1}$ from $\mathbf{Z}_{3}$, which would enable analysis of indirect and spurious discrimination. Future work could consider discovery for other fairness estimands, such as the natural effects (Pearl 2014) and path-specific effects (Avin, Shpitser, and Pearl 2005).
Acknowledgments
This research was supported by NSF 2212175; NIH RF1AG084178, R01AG076448, R01AG080624, R01AG076234, R01AG080991 and RF1AG072449; and the NSF Graduate Research Fellowship (J. Maasch).
Appendix
Appendix A Further Discussion on Direct Effect Estimation
A.1 Comparison of Direct Effect Measures
Here we explore the differences and similarities between the CDE and alternative direct effect measures, and provide justification for our use of the CDE in this work.
Alternative Direct Effect Measures
We elaborate on the comparison between the CDE and its most popular counterpart, the NDE (Pearl 2001). [^11] The Ctf-DE is an extension of the NDE; see Zhang and Bareinboim (2018). The CDE measures the expected change in outcome as the exposure changes when mediators $\mathbf{M}$ are uniformly fixed to $\mathbf{m}$, a constant value over the population (as simulated by conditioning on $\mathbf{M}$; Pearl 2012a). We express the CDE in $do$-notation as
$\mathrm{CDE}(\mathbf{m})$ $\coloneqq\mathbb{E}[Y_{x,\mathbf{m}}-Y_{x^{*},\mathbf{m}}]$ (7)
$=\mathbb{E}[Y|do(X=x,\mathbf{M}=\mathbf{m})]-\mathbb{E}[Y|do(X=x^{*},\mathbf{M}=\mathbf{m})].$ (8)
Assuming access to a covariate set $\mathbf{S}$ that simultaneously controls for confounding of the exposure-outcome and mediator-outcome relationships such that the conditions described in Definition 5 are satisfied, [^12] When a single set $\mathbf{S}$ is not sufficient for valid control of confounding for both $\{X,Y\}$ and $\{\mathbf{M},Y\}$, we can apply a more complex formula that requires two covariate sets. See Vanderweele 2011 (Proposition 1) and Pearl 2014 (Appendix C) for elaboration. the CDE is identifiable from observational data as
$\sum_{\mathbf{s}}\big{(}\mathbb{E}[Y|x,\mathbf{s},\mathbf{m}]-\mathbb{E}[Y|x^{*},\mathbf{s},\mathbf{m}]\big{)}P(\mathbf{s}).$ (9)
To obtain a unique estimate irrespective of potential interactions between $X$ and $\mathbf{M}$, we marginalize over the mediator values:
$\mathrm{WCDE}$ $\coloneqq\sum_{\mathbf{m}}\sum_{\mathbf{s}}\big{(}\mathbb{E}[Y|x,\mathbf{s},\mathbf{m}]-\mathbb{E}[Y|x^{*},\mathbf{s},\mathbf{m}]\big{)}P(\mathbf{s})P(\mathbf{m}).$ (10)
In contrast, the NDE measures the expected change in outcome as the exposure changes, with mediators fixed at the values that they would have been per individual prior to this change in exposure. The NDE captures the portion of the total causal effect that would be transmitted to the outcome if the mediators did not respond to changes in the exposure.
Definition A.1 (Natural direct effect (NDE), Pearl 2014).
We express the average NDE as
$\mathrm{NDE}(x,x^{*};Y)$ $=\mathbb{E}[Y_{x,\mathbf{M}_{x}^{*}}-Y_{x^{*},\mathbf{M}_{x}^{*}}].$ (11)
where $\mathbf{M}_{x}^{*}$ takes the value that the individual would have attained under $X=x^{*}$. The NDE requires probabilities of nested counterfactuals and a cross-world counterfactual assumption (Shpitser and VanderWeele 2011; Pearl 2012b, 2014; Andrews and Didelez 2021). Nevertheless, Pearl (2001) showed that the NDE can be expressed with do-notation and reduced to a do-free expression by valid covariate adjustment, allowing for identification and estimation (Pearl 2012b). The average NDE in Markovian models is identifiable from observational data per Corollary 2 in Pearl (2001), allowing counterfactual expressions to be replaced by probabilistic estimands:
$\mathrm{NDE}(x,x^{*};Y)$ $=\sum_{\mathbf{m}}\sum_{\mathbf{s}}\big{(}\mathbb{E}[Y|x,\mathbf{s},\mathbf{m}]-\mathbb{E}[Y|x^{*},\mathbf{s},\mathbf{m}]\big{)}P(\mathbf{s})P(\mathbf{m}|x^{*},\mathbf{s}).$ (12)
Note that the WCDE and NDE differ only by the way that we weigh the CDE: for the WCDE, we weigh by $P(\mathbf{m})$, and for the NDE, we weigh by $P(\mathbf{m}|x^{*},s)$. This stems from the fact that the CDE is a function of interventional distributions, while the NDE is a counterfactual quantity. Meanwhile, the Ctf-DE measures NDEs conditioned on the subpopulation $X=x$ (Zhang and Bareinboim 2018).
Total Effect Decomposition
Unlike the NDE, the CDE does not admit a valid decomposition of the total causal effect into direct and indirect effects for the general nonparametric setting (Pearl 2001). Conversely, the total effect is always equal to the difference of the NDE and the reverse transition of the natural indirect effect (NIE), even in the presence of nonlinearities and exposure-mediator interactions (Pearl 2014). However, in the absence of exposure-mediator interaction, the CDE and NDE are equivalent and the CDE can be used for total effect decomposition (Vanderweele 2011). Guaranteed total effect decomposability is a primary reason that the natural effects are employed in CFA (Plečko and Bareinboim 2024). However, when qualitative indicators of direct discrimination are all that is needed, it suffices to estimate the CDE (Zhang and Bareinboim 2018).
A.2 Graphical Criteria for Direct Effect Identification
Sufficient and necessary conditions for identification vary by estimand, but generally require covariate adjustment (VanderWeele 2019). Formal graphical rules can allow the VAS to be read directly from causal structures, even in the presence of latent confounding (Perkovic et al. 2015; Maathuis and Colombo 2015; Runge 2021). Graphical criteria for valid adjustment have been proposed for diverse causal quantities (e.g., Pearl’s backdoor criterion for total effect estimation; Pearl 1995). Often, multiple valid criteria exist per estimand. These vary in the amount of prior structural knowledge required, and criteria that require less knowledge are easier to implement in practice (VanderWeele and Shpitser 2013; Guo, Lundborg, and Zhao 2022).
Several graphical criteria have been proposed for the identification of direct effects and related quantities. Shpitser and Pearl (2006) present a complete graphical criterion for conditional interventional distributions. Pearl (2001) defines a graphical criterion for identifying the NDE from observational data (Corollary 1). Discussion, debate, and confusion around the sufficient and necessary conditions for NDE identification have inspired a flurry of papers (Imai, Keele, and Yamamoto 2010; Imai, Keele, and Tingley 2010), as summarized in Pearl (2014). Shpitser and VanderWeele (2011) present a complete graphical criterion for the NDE that reconciles some debate, showing where common criteria coincide.
Graphical criteria for direct effect estimation are often challenging for practitioners to implement (Shpitser and VanderWeele 2011). To increase the practicality of covariate adjustment in complex and low-knowledge domains, we advocate for the development of graphical criteria that are (1) relatively intuitive and practical to implement in applied research, and (2) are satisfied by knowledge that can be reliably learned through data-driven methods. Thus, we apply the causal partition taxonomy of Maasch et al. (2024) to provide local graphical intuition for fast and practical adjustment set discovery.
Appendix B Proofs
B.1 Proof of Theorem 1 (Correctness of Algorithm 1)
Proof.
We show that each set of tests is correct when all stated assumptions are met.
  • $\mathbf{Z}_{8}$ (Line 3): As proven in (Maasch et al. 2024), the independence relations tested in Line 3 uniquely define partition $\mathbf{Z}_{8}$ and are therefore an asymptotically correct test for identifying this partition, and this partition only.
  • $\mathbf{Z}_{5,7}$ (Line 4): As proven in (Maasch et al. 2024), the independence relations tested in Line 4 define partition $\mathbf{Z}_{7}$ in all structures and $\mathbf{Z}_{5}$ in structures for which $|\mathbf{Z}_{1}|=0$ (i.e., no confounders, latent nor observed, exist for $X$ and $Y$). Therefore, this test is asymptotically guaranteed to return all $\mathbf{Z}_{7}$ for arbitrary structures, and all $\mathbf{Z}_{5}$ if $\mathbf{Z}_{1}$ is the empty set.
  • $\mathbf{Z}_{4}$ (Line 5): As proven in (Maasch et al. 2024), the independence relations tested in Line 5 uniquely define partition $\mathbf{Z}_{4}$ and are therefore an asymptotically correct test for identifying this partition, and this partition only.
  • $\mathbf{Z}_{1\in pa(Y)}\cup\mathbf{Z}_{3\in pa(Y)}$ (Line 8): At this stage, the only partitions remaining under consideration are $\mathbf{Z}_{1}$, $\mathbf{Z}_{2\notin de(Y)}$, $\mathbf{Z}_{3}$, and $\mathbf{Z}_{5}$ (assuming that $\mathbf{Z}_{1}$ is not the empty set). By definition, no $\mathbf{Z}_{5}$ will ever be adjacent to $Y$, as this partition cannot share a causal path to $Y$ that is not mediated by $X$ (Hernán and Robins 2006; Lousdal 2018; Maasch et al. 2024). All $\mathbf{Z}_{2\notin de(Y)}$ are descended from $\mathbf{Z}_{1}$, $\mathbf{Z}_{3}$, and/or $\mathbf{Z}_{4}$ (i.e., any partition that can enter into $Y$), but not from $Y$ itself. Thus, all $\mathbf{Z}_{2}$ and $\mathbf{Z}_{5}$ will be eliminated by eliminating variables not adjacent to $Y$ given $X\cup\mathbf{Z}_{4}\cup\{\mathbf{Z}^{\prime}\setminus Z\}$. This test will also result in the elimination of $\mathbf{Z}_{1}$ and $\mathbf{Z}_{3}$ that are not adjacent to $Y$, including proxy confounders and proxy mediators as defined in Maasch et al. (2024).
  • $\mathbf{Z}_{4\in pa(Y)}$ (Line 10): Any intervening node on the causal path from $\mathbf{Z}_{4}$ to $Y$ ($\mathbf{Z}_{4}\to\cdots\to Y$) must also belong to a partition that can enter into $Y$ ($\cdots\to Y$). By definition, this is strictly $\mathbf{Z}_{1}$, $\mathbf{Z}_{3}$, and $\mathbf{Z}_{4}$. Therefore, to obtain $d$-separation between $Y$ and any member of $\mathbf{Z}_{4}$ that is not adjacent to $Y$, a correct conditioning set will contain $X$ and members of $\mathbf{Z}_{1}$, $\mathbf{Z}_{3}$ and $\mathbf{Z}_{4}$. Conditioning on $X$ is essential, as the remaining conditioning set can open noncausal paths between $\mathbf{Z}_{4}$ and $Y$. It is therefore sufficient to condition on all $\mathbf{Z}_{1}$ and $\mathbf{Z}_{3}$ that are already known to be adjacent to $Y$, unioned with $X$ and all known $\mathbf{Z}_{4}$. Any variable that is marginally dependent on $Y$ given this conditioning set is a member of $\mathbf{Z}_{4}$ that is adjacent to $Y$.
  • $\mathbf{A}_{\text{{DE}}}$ (Line 11): By definition of $\mathbf{A}_{\text{{DE}}}$.
  • SDC (Line 12): By definition of the SDC (Definition 3), $SDC=\mathbf{1}(X\in pa(Y))$. If $X$ is not a parent of $Y$, it will always be $d$-separable from $Y$ when all their confounding and mediating paths are blocked. If $X$ is a parent of $Y$, $X$ and $Y$ will be always conditionally dependent given any conditioning set. Therefore, the SDC will evaluate to 0 if and only if $X\perp\!\!\!\perp Y|\mathbf{Z}_{1\in pa(Y)}\cup\mathbf{Z}_{3\in pa(Y)}$.
B.2 Proof of Theorem 2
We show that latent variables that are not parents of $Y$ do not impact the asymptotic guarantees of Algorithm 1.
Proof.
Maasch et al. (2024) have shown previously that the marginal and conditional independence tests used to identify $\mathbf{Z}_{4}$ and $\mathbf{Z}_{8}$ are not affected by latent variables. Intuition for this follows from the fact that these tests only require knowledge of $\{X,Y\}$ and the candidate variable $Z$. In Algorithm 1, the test for $\mathbf{Z}_{5,7}$ also relies only on knowledge of $X$, $Y$, and $Z$, and is therefore correct in the presence of latent variables. As we do not require identification of $\mathbf{Z}_{5}$ itself for downstream results, this test suffices. Otherwise, latent variables could pose identifiability problems for $\mathbf{Z}_{5}$, as in prior works (Maasch et al. 2024). A2 alone is sufficient to ensure $d$-separation between $Y$ and its non-parents at Lines 8 and 10, guaranteeing identification of the true parents of $Y$ in $\mathbf{Z}_{1}$, $\mathbf{Z}_{3}$, and $\mathbf{Z}_{4}$. Therefore, latent variables that are not parents of $Y$ have no impact on Algorithm 1. ∎
B.3 Proof of Theorem 3
Proof.
This theorem claims that $\mathbf{A}_{\text{{DE}}}\coloneqq\mathbf{Z}_{1\in pa(Y)}\cup\mathbf{Z}_{3\in pa(Y)}\cup\mathbf{Z}_{4\in pa(Y)}$ is a VAS for WCDE identification under the conditions described in Section 2.2. Per the identifiability conditions of the WCDE (Definition 5), a VAS must block all backdoor paths for $\{X,Y\}$ and $\{\mathbf{M},Y\}$. As shown in Pearl (2001), knowledge of all mediators for $\{X,Y\}$ that are parents of $Y$ is sufficient for identifying the CDE (i.e., knowledge of every mediator in the full causal graph is unnecessary). Thus, we show that adjusting for $\mathbf{A}_{\text{{DE}}}$ (1) blocks all backdoor paths for $\{X,Y\}$, (2) blocks all backdoor paths for $\{\mathbf{M},Y\}$, (3) blocks all frontdoor paths for $\{X,Y\}$, and (4) does not open non-causal paths by conditioning on colliders. We use the language of causal partitions defined in Maasch et al. (2024).
  1. 1.
    Backdoor paths for $\{X,Y\}$ are blocked. Any confounder for $\{X,Y\}$ will be in causal partition $\mathbf{Z}_{1}$, by definition of $\mathbf{Z}_{1}$ (Table 1; Maasch et al. 2024). All backdoor paths enter $Y$ by definition (i.e., $\cdots\to Y$). As $\mathbf{A}_{\text{{DE}}}$ contains all confounders that are parents of $Y$, all backdoor paths for $\{X,Y\}$ will be blocked when conditioning on $\mathbf{A}_{\text{{DE}}}$.
  2. 2.
    Backdoor paths for $\{\mathbf{M},Y\}$ are blocked. Any confounder for $\mathbf{M}$ and $Y$ will be in $\mathbf{Z}_{1}\cup\mathbf{Z}_{3}\cup\mathbf{Z}_{4}$ (e.g., Figure B.1), as these are the only causal partitions with edges pointing into $Y$. Set $\mathbf{A}_{\text{{DE}}}$ by definition contains all members of $\mathbf{Z}_{1}$, $\mathbf{Z}_{3}$, and $\mathbf{Z}_{4}$ that are parents of $Y$.
  3. 3.
    Frontdoor paths for $\{X,Y\}$ are blocked. Any mediator for $\{X,Y\}$ will be in causal partition $\mathbf{Z}_{3}$, by definition of $\mathbf{Z}_{3}$ (Table 1; Maasch et al. 2024). All frontdoor paths enter $Y$ by definition (i.e., $\cdots\to Y$). As $\mathbf{A}_{\text{{DE}}}$ contains all mediators that are parents of $Y$, all frontdoor paths will be blocked.
  4. 4.
    Non-causal paths through colliders are blocked. Though $\mathbf{Z}$ can contain colliders in $\mathbf{Z}_{2\notin de(Y)}$, these cannot be parents of $Y$ by definition. As defined, $\mathbf{A}_{\text{{DE}}}$ does not contain any members of $\mathbf{Z}_{2}$. Thus, non-causal paths will not be opened by conditioning on $\mathbf{A}_{\text{{DE}}}$.
[…]
Figure B.1: $Y$ and members of $\mathbf{Z}_{3}$ can be confounded by members of $\mathbf{Z}_{1}$ (e.g., left), $\mathbf{Z}_{3}$ (e.g., center), or $\mathbf{Z}_{4}$ (e.g., right), as these are the only causal partitions with edges pointing into $Y$.
Appendix C Beyond Bias: Variance and Actionability
Definition 8 defines a VAS for the WCDE (Equation 5) in the general setting, irrespective of A1 and A2. However, in settings that assume A1 and A2, bypassing structure learning and adjusting for all $\mathbf{Z}$ does not induce bias in WCDE estimation. However, this is not advised for two reasons: (1) this risks unnecessary adjustment, which can inflate the variance of causal effect estimators under finite data (Schisterman, Cole, and Platt 2009) and (2) ignoring causal structure limits the informativeness and actionability of fairness conclusions.
Impacts on Asymptotic Variance
While $\mathbf{Z}$ itself constitutes a VAS under Assumptions A1 and A2, adjusting for all $\mathbf{Z}$ with no knowledge of its causal structure can undermine statistical efficiency. We highlight some prior theorems here, using the language of causal partitions.
Consider the optimality of a VAS with respect to the asymptotic variance (AV) of the causal effect estimator. For total effect estimation in both linear and nonparametric settings, it has been proven that the asymptotically optimal VAS captures less or equal information about the exposure and more or equal information about the outcome than alternative VAS (Rotnitzky and Smucler 2020; Witte et al. 2020; Henckel, Perković, and Maathuis 2022). This implies: [^13] See Theorem 1 and its corollaries in Henckel, Perković, and Maathuis (2022), which were later proven for the nonparametric case in Rotnitzky and Smucler (2020).
  1. 1.
    VAS containing members of $\mathbf{Z}_{5}$ will have greater or equal AV than VAS not containing any $\mathbf{Z}_{5}$;
  2. 2.
    VAS containing members of $\mathbf{Z}_{4}$ will have lesser or equal AV than VAS not containing any $\mathbf{Z}_{4}$; and
  3. 3.
    VAS containing $\mathbf{Z}_{1\not\in pa(Y)}$ or $\mathbf{Z}_{4\not\in pa(Y)}$ will have greater or equal AV than VAS where all $\mathbf{Z}_{1}$ and $\mathbf{Z}_{4}$ are in $pa(Y)$.
While these results apply to total effect estimation, we follow this intuition for WCDE estimation using $\mathbf{A}_{\text{{DE}}}$, which excludes $\mathbf{Z}_{5}$ and $\mathbf{Z}_{1,4\not\in pa(Y)}$. We defer formal derivation of the asymptotically optimal VAS for the WCDE to future work. In Section 5 and Appendix D.5, we empirically demonstrate variance inflation in finite samples when adjusting for all $\mathbf{Z}$ versus $\mathbf{A}_{\text{{DE}}}$.
Actionable Fairness Conclusions
Ultimately, CFA must support our ability to take actions that remove unwanted association between the protected attribute and outcome. In CFA, we examine causal structure to (1) determine whether unfairness is present and (2) identify structural mechanisms that can be redressed through interventions (e.g., policy change). LD3 addresses goal (1) by using the SDC and WCDE for binary classification of the SCM (fair/unfair) with respect to direct discrimination. Through causal partition labeling, LD3 yields interpretable VAS that provide partial information for goal (2). Adjusting for all $\mathbf{Z}$ in WCDE estimation with no prior structure learning can address (1), but not (2). Consider the following example SCMs with unknown graphical structure, where A1 and A2 hold.
  1. 1.
    (A) WCDE $\neq$ 0, (B) $\mathbf{Z}_{1,3}$ is empty. We can assess (A) by adjusting for $\mathbf{Z}$ without structure learning, but we cannot know (B). Thus, we can conclude that the SCM is unfair with respect to direct discrimination, but not that indirect and spurious unfairness are absent. Discovery methods like LD3 address (A) and (B) by enabling precise WCDE estimation and indicating that policy interventions need only address direct mechanisms of unfairness.
  2. 2.
    (A) WCDE = 0, (B) $\mathbf{Z}_{1,3}$ is not empty. All pathways of unfairness in the SCM pass through $\mathbf{Z}_{1,3\in pa(Y)}$. We can ask the policy design question: are any variables in $\mathbf{Z}_{1,3\in pa(Y)}$ feasible points of intervention for mitigating unfairness? While (A) implies (B) even without structure learning, we cannot answer our policy design question without the structural knowledge returned by methods like LD3.
Empirical demonstration of improved VAS interpretability is given in Section 5. Examples of structural insights provided by local discovery are given in the real-world CFA described in Section 6.
Appendix D Empirical Validation on Synthetic and Semi-Synthetic Data
Computing Resources
All experiments used an Apple MacBook (M2 Pro Chip; 12 CPU cores; 16G memory).
D.1 LD3 in the Presence of Latent Variables
We explored the impacts of latent variables that do and do not violate Assumption A2 (Table D.1). We iteratively dropped individual variables from the observed data represented by Figure D.1, including variables that do and do not act as confounders.
Results illustrate that (1) Theorem 2 holds and (2) Assumption A2 is sufficient but not necessary. WCDE for latent confounders $B_{1}$, $M_{1}$, and $Z_{4a}$ remain unbiased, as these are not adjacent to $Y$ (Theorem 2). Though latent $M_{2}$ could induce bias by allowing for the erroneous inclusion of $M_{3}$ in $\mathbf{A}_{\text{{DE}}}$, WCDE results are not significantly impacted in this setting. Additionally, though $Z_{3c}$ is a parent of $Y$, dropping this variable simply causes LD3 to place its parent $Z_{3b}$ in $\mathbf{A}_{\text{{DE}}}$. In this setting, the predicted $\mathbf{A}_{\text{{DE}}}$ has 100% F1, precision, and recall with respect to the expected VAS, and the WCDE remains unbiased ($1.25[1.24,1.26]$). Thus, Assumption A2 is sufficient but not necessary for returning a VAS.
Latent WCDE SDC Acc $\mathbf{A}_{\text{{DE}}}$ F1 $\mathbf{A}_{\text{{DE}}}$ Prec $\mathbf{A}_{\text{{DE}}}$ Rec
$Z_{1}\in\mathbf{Z}_{1}$ 1.70 [1.66,1.75] 1.0 0.80 [0.80,0.80] 0.67 [0.67,0.67] 1.00 [1.00,1.00]
$B_{1}\in\mathbf{Z}_{1}$ 1.25 [1.24,1.26] 1.0 1.00 [1.00,1.00] 1.00 [1.00,1.00] 1.00 [1.00,1.00]
$B_{2}\in\mathbf{Z}_{1}$ 1.22 [1.17,1.27] 1.0 0.91 [0.89,0.93] 0.84 [0.81,0.86] 1.00 [1.00,1.00]
$B_{3}\in\mathbf{Z}_{1}$ 1.72 [1.68,1.76] 1.0 0.82 [0.8,0.840] 0.70 [0.67,0.73] 1.00 [1.00,1.00]
$M_{1}\in\mathbf{Z}_{5}$ 1.25 [1.24,1.26] 1.0 0.99 [0.98,1.01] 0.99 [0.96,1.01] 1.00 [1.00,1.00]
$M_{2}\in\mathbf{Z}_{4}$ 1.26 [1.25,1.26] 1.0 0.86 [0.86,0.86] 0.75 [0.75,0.75] 1.00 [1.00,1.00]
$Z_{4a}\in\mathbf{Z}_{4}$ 1.25 [1.24,1.26] 1.0 1.00 [1.00,1.00] 1.00 [1.00,1.00] 1.00 [1.00,1.00]
Table D.1: Results with latent confounders. WCDE estimates (WCDE), SDC accuracy (Acc), $\mathbf{A}_{\text{{DE}}}$ F1, $\mathbf{A}_{\text{{DE}}}$ precision (Prec), and $\mathbf{A}_{\text{{DE}}}$ recall (Rec) for LD3 as confounders in Figure D.1 are iteratively dropped from the observed data (linear-Gaussian data; Fisher-z tests; $\alpha=0.01$; true direct effect = 1.25). Values are means over 10 replicates with 95% confidence intervals in brackets. Latent nodes in red are adjacent to outcome $Y$, violating Assumption A2. $\mathbf{A}_{\text{{DE}}}$ metrics were scored by comparing the set returned by LD3 to the true $\mathbf{A}_{\text{{DE}}}$ set-minus the latent variable.
[…]
Figure D.1: Eighteen-node DAG used to evaluate the ability of LD3 to recover the true $\mathbf{A}_{\text{{DE}}}$, assess the SDC, and estimate the weighted CDE. For WCDE experiments (Figures D.3, D.4), all nodes were observed. For latent variable experiments (Section D.1), one dashed node was treated as latent per iteration.
D.2 Baseline Evaluation
Baseline Evaluation for Parent Discovery
Baselines were evaluated on the basis of F1, precision, and recall with respect to the true parents of outcome $Y$ (i.e., the members of $\mathbf{A}_{\text{{DE}}}$). Only LD3, LDECC, and MB-by-MB directly return the parents of $Y$. Additionally, PC, LDECC, and MB-by-MB return results relative to the MEC. Therefore, we evaluate each baseline with at least one form of post-processing to adjust for the task of $\mathbf{A}_{\text{{DE}}}$ prediction. Post-processing is described below.
  1. 1.
    NOTEARS: Retain all parents of the $Y$ in the unique graph.
  2. 2.
    DirectLiNGAM: Retain all parents of the $Y$ in the unique graph.
  3. 3.
    PC: Retain the intersection of parent sets across DAGs in the MEC.
  4. 4.
    PC: Retain the union of parent sets across DAGs in the MEC.
  5. 5.
    MB-by-MB: Retain the variables labeled as parents of the target.
  6. 6.
    LDECC: Retain the variables labeled as parents of the target.
  7. 7.
    LDECC: Retain the variables labeled as parents of the target unioned with the variables adjacent to $Y$ that are unoriented in the MEC.
  8. 8.
    LDP: Retain $\mathbf{Z}_{1}\cup\mathbf{Z}_{\textsc{POST}}\cup\mathbf{Z}_{4}$. As our experimental setting requires that there are no descendants of the outcome, the superset $\mathbf{Z}_{\textsc{POST}}\coloneqq\mathbf{Z}_{2\in de(Y)}\cup\mathbf{Z}_{3}\cup\mathbf{Z}_{6}$ defined in Maasch et al. (2024) will now only contain $\mathbf{Z}_{3}$.
  9. 9.
    LDPpa: Retain the variables that are not conditionally independent of $Y$ given the remaining members of the set.
For the Sangiovese benchmark, we compared LD3 to NOTEARS, DirectLiNGAM, MB-by-MB, LDECC, PC, and LDPpa. For Asia and Sachs, we compared LD3 to PC, PC, LDECC, LDECC, LDP with no post-processing, and LDPpa to emphasize the impacts of different post-processing methods.
D.3 Erdős-Rényi Graphs
Figure D.2: Total runtime and total CI tests performed with respect to total edges and total nodes, using an oracle on random directed acyclic Erdős-Rényi graphs (Erdős, Rényi et al. 1960).
D.4 Weighted CDE Estimation
Results
To evaluate SDC and WCDE correctness, we tested LD3 on linear-Gaussian instantiations of an 18-node DAG with an M-structure (i.e., a member of $\mathbf{Z}_{2\notin de(Y)}$; Ding and Miratrix 2015) and multiple members of $\mathbf{Z}_{1}$, $\mathbf{Z}_{3}$, and $\mathbf{Z}_{4}$ (Figure D.1). The WCDE was estimated using linear regression with $\mathbf{A}_{\text{{DE}}}$ as covariates. $\mathbf{A}_{\text{{DE}}}$ F1, precision, and recall and SDC accuracy converged toward 100% as sample size increased. The WCDE converged toward the true direct effect with low variance (Figures D.3, D.4).
Figure D.3: Mean WCDE estimates over 100 replicates per sample size for linear-Gaussian instantiations of Figure D.1 (95% confidence intervals in shaded regions). Ground truth WCDE is denoted by the dashed line. LD3 used parametric Fisher-z tests ($\alpha=0.01$), as the data generating process was known for this simulation. Estimates were obtained with linear regression (https://scikit-learn.org/). Note: This pipeline is for ease of explication in the illustrative setting only. For unknown data generating processes, we do not recommend this pipeline due to restrictive parametric assumptions.
Figure D.4: SDC accuracy, $\mathbf{A}_{\text{{DE}}}$ F1, precision, and recall across sample sizes when SDC = 0 and SDC = 1. Values were averaged over 100 replicates of linear-Gaussian instantiations of Figure D.1, with 95% confidence intervals in shaded regions. LD3 used parametric Fisher-z tests ($\alpha=0.01$), as the data generating process was known for this simulation.
D.5 Unnecessary Adjustment: Estimator Variance and VAS Interpretability
[…]
Figure D.5: The true $\mathbf{A}_{\text{{DE}}}$ for $X$ and $Y$ is in gray.
$\mathbf{A}_{\text{{DE}}}$ All $\mathbf{Z}$
$n$ Mean Variance Mean Variance
100 6.98 0.83 6.97 7.79
1000 7.17 0.86 7.28 8.27
10000 6.91 1.16 7.25 9.04
Table D.2: Given the DAG in Figure D.5, adjusting for $\mathbf{Z}$ inflates WCDE variance relative to $\mathbf{A}_{\text{{DE}}}$ in finite samples (size $n$). True WCDE was 7. Data generating process was linear-Gaussian. WCDE estimation used linear regression with 100 replicates per sample size.
All $\mathbf{Z}$ True $\mathbf{A}_{\text{{DE}}}$ Pred $\mathbf{A}_{\text{{DE}}}$
$n$ Mean Variance Mean Variance Mean Variance $\mathbf{A}_{\text{{DE}}}$ F1
500 0.239 0.052 0.347 0.004 0.344 0.004 0.99 [0.98,1.0]
1000 -0.011 0.038 0.35 0.003 0.349 0.003 0.99 [0.98,1.0]
10000 0.151 0.013 0.345 0.000 0.344 0.000 0.99 [0.98,1.0]
Table D.3: Statistical Efficiency. WCDE estimate mean and variance when adjusting for all $\mathbf{Z}$, the true $\mathbf{A}_{\text{{DE}}}$ defined by our graphical criterion (True $\mathbf{A}_{\text{{DE}}}$), and the predicted $\mathbf{A}_{\text{{DE}}}$ returned by LD3 (Pred $\mathbf{A}_{\text{{DE}}}$) for the DAG in Figure D.5. Variance using all $\mathbf{Z}$ was at least 12.6×12.6$\times$12.6 × higher than using Pred $\mathbf{A}_{\text{{DE}}}$. F1 was high and variance for Pred $\mathbf{A}_{\text{{DE}}}$ was equal to the variance using the ground truth $\mathbf{A}_{\text{{DE}}}$. Variables were categorical with quadratic causal functions. WCDE estimation used double machine learning using econml (https://econml.azurewebsites.net/). Means are over 100 replicates per sample size (95% confidence intervals in brackets). Though adjusting for all $\mathbf{Z}$ will not theoretically induce bias in this setting, we see that mean WCDE values do diverge (and sometimes flip sign).
$n$ $|\mathbf{Z}|$ $|\mathbf{A}_{\text{{DE}}}|$ Acc F1 Rec Prec Time Ecoli $1k$ 44 2.0 [2.0,2.0] 1.00 [1.00,1.00] 1.00 [1.00,1.00] 1.00 [1.00,1.00] 1.00 [1.00,1.00] 0.016 [0.01,0.021] Andes $50k$ 221 4.5 [4.06,4.94] 1.00 [1.00,1.00] 0.95 [0.9,0.99] 0.91 [0.83,0.99] 1.00 [1.00,1.00] 0.76 [0.737,0.784]
Table D.4: Interpretability. $\mathbf{A}_{\text{{DE}}}$ discovered by LD3 can provide greater interpretability than adjusting for all $\mathbf{Z}$, especially in large DAGs. We used the binary Andes (Conati et al. 1997) and Gaussian Ecoli (Schäfer and Strimmer 2005) benchmarks from bnlearn (Scutari 2010). When the true parent set is small but total number of nodes is moderate (e.g., Ecoli) to large (e.g., Andes), forgoing structure learning by adjusting for $\mathbf{Z}$ is a missed opportunity to provide mechanistic insights (e.g., causal fairness mechanisms). In the fairness setting, we can draw direct discrimination conclusions about the members of $\mathbf{A}_{\text{{DE}}}$, but not $\mathbf{Z}$ alone. On average, eliminating extraneous variables took less than one second. Ecoli used Fisher-z tests ($\alpha=0.001$) and Andes used $\chi^{2}$ ($\alpha=0.01$). Values were averaged over ten replicates (95% confidence intervals in brackets).
D.6 bnlearn Benchmarks
$n$ Method Time CI tests Accuracy F1 Precision Recall 250 LD3 0.005 [0.003,0.008] 60.8 [60.41,61.19] 0.74 [0.71,0.78] 0.7 [0.65,0.76] 1.00 [1.00,1.00] 0.55 [0.48,0.62] LDECC 0.092 [0.063,0.12] 316.6 [220.98,412.22] 0.61 [0.55,0.67] 0.47 [0.33,0.6] 0.9 [0.7,1.0] 0.32 [0.2,0.43] MB-by-MB 0.092 [0.047,0.136] 268.4 [179.54,357.26] 0.63 [0.57,0.68] 0.5 [0.37,0.62] 0.9 [0.7,1.0] 0.35 [0.25,0.45] LDP 0.009 [0.006,0.012] 112.1 [90.82,133.38] 0.76 [0.72,0.81] 0.76 [0.71,0.8] 0.94 [0.87,1.0] 0.64 [0.59,0.68] PC 0.118 [0.099,0.136] 557.9 [524.73,591.07] 0.64 [0.58,0.69] 0.51 [0.39,0.64] 0.9 [0.7,1.0] 0.36 [0.27,0.46] DirectLiNGAM 0.142 [0.138,0.145] - 0.63 [0.55,0.7] 0.56 [0.43,0.69] 0.85 [0.75,0.95] 0.48 [0.32,0.63] NOTEARS 1.216 [0.985,1.447] - 0.79 [0.75,0.82] 0.79 [0.75,0.82] 0.91 [0.86,0.96] 0.7 [0.65,0.75] 500 LD3 0.006 [0.004,0.008] 60.8 [60.41,61.19] 0.85 [0.79,0.91] 0.84 [0.77,0.91] 1.00 [1.00,1.00] 0.74 [0.64,0.84] LDECC 0.137 [0.104,0.169] 458.2 [354.85,561.55] 0.68 [0.63,0.73] 0.6 [0.52,0.69] 0.98 [0.94,1.0] 0.45 [0.37,0.53] MB-by-MB 0.127 [0.113,0.141] 370 [328.41,411.59] 0.64 [0.61,0.68] 0.55 [0.48,0.61] 0.98 [0.94,1.0] 0.39 [0.32,0.46] LDP 0.01 [0.005,0.016] 84.3 [66.81,101.79] 0.78 [0.73,0.82] 0.77 [0.72,0.82] 0.94 [0.88,1.0] 0.66 [0.6,0.73] PC 0.181 [0.162,0.201] 879.6 [843.24,915.96] 0.61 [0.56,0.67] 0.47 [0.38,0.57] 1.00 [1.00,1.00] 0.32 [0.23,0.42] DirectLiNGAM 0.152 [0.144,0.16] - 0.56 [0.46,0.67] 0.42 [0.18,0.65] 0.43 [0.19,0.67] 0.41 [0.18,0.65] NOTEARS 1.092 [1.025,1.158] - 0.79 [0.76,0.82] 0.78 [0.75,0.81] 0.95 [0.9,1.0] 0.66 [0.63,0.7] 1000 LD3 0.009 [0.007,0.011] 61 [61,61] 0.97 [0.95,0.99] 0.99 [0.95,0.99] 1.00 [1.00,1.00] 0.95 [0.91,0.99] LDECC 0.217 [0.185,0.248] 669.3 [560.92,777.68] 0.66 [0.61,0.72] 0.57 [0.45,0.68] 1.00 [1.00,1.00] 0.41 [0.31,0.52] MB-by-MB 0.201 [0.175,0.228] 592.7 [538.57,646.83] 0.69 [0.66,0.72] 0.63 [0.58,0.68] 1.00 [1.00,1.00] 0.46 [0.41,0.51] LDP 0.006 [0.003,0.009] 75 [63.46,86.54] 0.95 [0.91,0.99] 0.95 [0.91,0.99] 0.99 [0.96,1.0] 0.92 [0.87,0.98] PC 0.289 [0.272,0.306] 1333.9 [1249.08,1418.72] 0.55 [0.5,0.6] 0.33 [0.21,0.45] 0.8 [0.54,1.06] 0.21 [0.13,0.29] DirectLiNGAM 0.366 [0.326,0.407] - 0.62 [0.55,0.69] 0.56 [0.4,0.73] 0.66 [0.51,0.8] 0.54 [0.34,0.73] NOTEARS 1.694 [1.571,1.818] - 0.8 [0.77,0.83] 0.79 [0.75,0.82] 1.00 [1.00,1.00] 0.65 [0.6,0.7]
Table D.5: Results on the Sangiovese benchmark from bnlearn (Scutari 2010), a linear-Gaussian model of Tuscan Sangiovese grape production (Magrini, Di Blasi, and Stefanini 2017). Ten replicate datasets were sampled at $n=[250,500,1000]$. Means are given with 95% confidence intervals in brackets. Exposure was pH and outcome was GrapeW. All constraint-based methods used Fisher-z tests ($\alpha=0.01$). NOTEARS hyperparameters were $\lambda_{1}=0.001$ and loss type $l_{2}$. DirectLiNGAM: https://lingam.readthedocs.io/; PC, LDECC, MB-by-MB: https://github.com/acmi-lab/local-causal-discovery; NOTEARS: https://github.com/xunzheng/notears; LDP: https://github.com/jmaasch/ldp.
Figure D.6: Total number of CI tests performed for the Asia (Lauritzen and Spiegelhalter 1988) and Sachs (Sachs et al. 2005) benchmarks from bnlearn (https://www.bnlearn.com/bnrepository/) (Scutari 2010) ($\chi^{2}$ tests; $\alpha=0.001$; 95% confidence intervals in shaded regions). For Asia (either $\to$ dysp), $|\mathbf{Z}|=6$ and $|\mathbf{A}_{\text{{DE}}}|=1$. For Sachs (erk $\to$ akt), $|\mathbf{Z}|=11$ and $|\mathbf{A}_{\text{{DE}}}|=1$. Raw data are reported in Tables D.7 and D.9.
Asia (Lung Cancer) SDC = 1: either $\to$ dysp SDC = 0: xray ↛$\not\to$ dysp F1 Prec Rec F1 Prec Rec Oracle LD3 1.00 1.00 1.00 1.00 1.00 1.00 PC 1.00 1.00 1.00 1.00 1.00 1.00 PC 1.00 1.00 1.00 1.00 1.00 1.00 LDECC 1.00 1.00 1.00 1.00 1.00 1.00 LDECC 1.00 1.00 1.00 1.00 1.00 1.00 LDPpa 1.00 1.00 1.00 1.00 1.00 1.00 LDP 0.50 0.33 1.00 0.67 0.50 1.00 $n=2.5k$ LD3 1.00 [1.00,1.00] 1.00 [1.00,1.00] 1.00 [1.00,1.00] 0.80 [0.69,0.91] 1.00 [1.00,1.00] 0.70 [0.54,0.86] PC 0.00 [0.00,0.00] 0.00 [0.00,0.00] 0.00 [0.00,0.00] 0.00 [0.00,0.00] 0.00 [0.00,0.00] 0.00 [0.00,0.00] PC 0.00 [0.00,0.00] 0.00 [0.00,0.00] 0.00 [0.00,0.00] 0.00 [0.00,0.00] 0.00 [0.00,0.00] 0.00 [0.00,0.00] LDECC 0.75 [0.61,0.89] 0.65 [0.46,0.84] 1.00 [1.00,1.00] 0.67 [0.49,0.85] 0.65 [0.46,0.84] 0.70 [0.54,0.86] LDECC 0.75 [0.61,0.89] 0.65 [0.46,0.84] 1.00 [1.00,1.00] 0.67 [0.49,0.85] 0.65 [0.46,0.84] 0.70 [0.54,0.86] LDPpa 1.00 [1.00,1.00] 1.00 [1.00,1.00] 1.00 [1.00,1.00] 0.83 [0.72,0.94] 1.00 [1.00,1.00] 0.75 [0.59,0.91] LDP 0.26 [0.24,0.28] 0.15 [0.14,0.16] 0.97 [0.90,1.00] 0.14 [0.13,0.14] 0.07 [0.07,0.08] 1.00 [1.00,1.00] $n=5k$ LD3 1.00 [1.00,1.00] 1.00 [1.00,1.00] 1.00 [1.00,1.00] 0.90 [0.80,1.00] 1.00 [1.00,1.00] 0.85 [0.70,1.00] PC 0.00 [0.00,0.00] 0.00 [0.00,0.00] 0.00 [0.00,0.00] 0.00 [0.00,0.00] 0.00 [0.00,0.00] 0.00 [0.00,0.00] PC 0.00 [0.00,0.00] 0.00 [0.00,0.00] 0.00 [0.00,0.00] 0.00 [0.00,0.00] 0.00 [0.00,0.00] 0.00 [0.00,0.00] LDECC 0.60 [0.47,0.73] 0.47 [0.29,0.64] 1.00 [1.00,1.00] 0.52 [0.36,0.68] 0.47 [0.29,0.64] 0.60 [0.47,0.73] LDECC 0.60 [0.47,0.73] 0.47 [0.29,0.64] 1.00 [1.00,1.00] 0.52 [0.36,0.68] 0.47 [0.29,0.64] 0.60 [0.47,0.73] LDPpa 1.00 [1.00,1.00] 1.00 [1.00,1.00] 1.00 [1.00,1.00] 0.93 [0.85,1.00] 1.00 [1.00,1.00] 0.90 [0.77,1.00] LDP 0.65 [0.62,0.68] 0.48 [0.45,0.52] 1.00 [1.00,1.00] 0.65 [0.62,0.67] 0.48 [0.45,0.51] 1.00 [1.00,1.00] $n=10k$ LD3 1.00 [1.00,1.00] 1.00 [1.00,1.00] 1.00 [1.00,1.00] 0.90 [0.80,1.00] 1.00 [1.00,1.00] 0.85 [0.70,1.00] PC 0.00 [0.00,0.00] 0.00 [0.00,0.00] 0.00 [0.00,0.00] 0.00 [0.00,0.00] 0.00 [0.00,0.00] 0.00 [0.00,0.00] PC 0.00 [0.00,0.00] 0.00 [0.00,0.00] 0.00 [0.00,0.00] 0.00 [0.00,0.00] 0.00 [0.00,0.00] 0.00 [0.00,0.00] LDECC 0.50 [0.50,0.50] 0.33 [0.33,0.33] 0.00 [0.00,0.00] 0.40 [0.40,0.40] 0.33 [0.33,0.33] 0.50 [0.50,0.50] LDECC 0.50 [0.50,0.50] 0.33 [0.33,0.33] 0.00 [0.00,0.00] 0.40 [0.40,0.40] 0.33 [0.33,0.33] 0.50 [0.50,0.50] LDPpa 1.00 [1.00,1.00] 1.00 [1.00,1.00] 1.00 [1.00,1.00] 0.90 [0.80,1.00] 1.00 [1.00,1.00] 0.85 [0.70,1.00] LDP 0.67 [0.67,0.67] 0.50 [0.50,0.50] 1.00 [1.00,1.00] 0.64 [0.61,0.67] 0.47 [0.44,0.5] 1.00 [1.00,1.00]
Table D.6: Empirical performance on the Asia (Lung Cancer) Bayesian network dataset (Lauritzen and Spiegelhalter 1988) from bnlearn (Scutari 2010). Binary data were evaluated with $\chi^{2}$ independence tests ($\alpha=0.001$; 95% confidence intervals in brackets). Unexpected decreases in baseline performance for increasing sample size might be due to the difficulty of this dataset for structure learning, as noted in the bnlearn documentation: “Standard learning algorithms are not able to recover the true structure of the network because of the presence of a node (E) with conditional probabilities equal to both 0 and 1” (https://www.bnlearn.com/documentation/man/asia.html).
Asia (Lung Cancer) SDC = 1: either $\to$ dysp SDC = 0: xray ↛$\not\to$ dysp Time Tests Time Tests Oracle LD3 0.002 29 0.002 31 PC 0.007 110 0.007 115 LDECC 0.003 25 0.004 25 LDP 0.003 43 0.003 50 $n=2.5k$ LD3 0.003 [0.002,0.003] 23.4 [22.56,24.24] 0.003 [0.003,0.003] 27.2 [26.56,27.84] PC 0.032 [0.03,0.034] 67.1 [63.31,70.89] 0.033 [0.031,0.035] 70.5 [66.79,74.21] LDECC 0.024 [0.016,0.033] 35.5 [23.78,47.22] 0.018 [0.013,0.023] 35.5 [23.78,47.22] LDP 0.003 [0.003,0.003] 27.6 [24.41,30.79] 0.003 [0.003,0.003] 31.0 [27.94,34.06] $n=5k$ LD3 0.004 [0.003,0.004] 22.9 [21.96,23.84] 0.004 [0.003,0.004] 26.6 [26.0,27.2] PC 0.056 [0.054,0.057] 73.6 [72.02,75.18] 0.058 [0.057,0.06] 77.1 [75.29,78.91] LDECC 0.049 [0.044,0.054] 63.5 [56.21,70.79] 0.047 [0.042,0.053] 63.5 [56.21,70.79] LDP 0.073 [0.056,0.091] 214.4 [175.9,252.9] 0.039 [0.033,0.044] 192.0 [167.74,216.26] $n=10k$ LD3 0.004 [0.004,0.004] 22.0 [22.0,22.0] 0.005 [0.005,0.005] 26.6 [26.0,27.2] PC 0.100 [0.097,0.103] 77.4 [75.9,78.9] 0.105 [0.102,0.109] 82.3 [80.38,84.22] LDECC 0.096 [0.083,0.109] 67.1 [65.34,68.86] 0.087 [0.084,0.09] 67.1 [65.34,68.86] LDP 0.073 [0.056,0.091] 214.4 [175.9,252.9] 0.039 [0.033,0.044] 192.0 [167.74,216.26]
Table D.7: Total number of independence tests performed and runtimes in seconds (95% confidence intervals in brackets).
Sachs Protein Expression SDC = 1: ERK $\to$ AKT SDC = 0: JNK ↛$\not\to$ P38 F1 Prec Rec F1 Prec Rec Oracle LD3 1.00 1.00 1.00 1.00 1.00 1.00 PC 1.00 1.00 1.00 1.00 1.00 1.00 PC 0.00 0.00 0.00 0.00 0.00 0.00 LDECC 1.00 1.00 1.00 1.00 1.00 1.00 LDECC 0.00 0.00 0.00 0.00 0.00 0.00 LDPpa 1.00 1.00 1.00 1.00 1.00 1.00 LDP 0.29 0.16 1.00 0.50 0.33 1.00 $n=5k$ LD3 0.97 [0.90,1.00] 0.95 [0.85,1.00] 1.00 [1.00,1.00] 0.96 [0.91,1.00] 0.93 [0.85,1.00] 1.00 [1.00,1.00] PC 0.00 [0.00,0.00] 0.00 [0.00,0.00] 0.00 [0.00,0.00] 0.77 [0.51,1.00] 0.80 [0.54,1.00] 0.75 [0.49,1.00] PC 0.00 [0.00,0.00] 0.00 [0.00,0.00] 0.00 [0.00,0.00] 0.73 [0.48,0.99] 0.80 [0.54,1.00] 0.70 [0.44,0.96] LDECC 1.00 [1.00,1.00] 1.00 [1.00,1.00] 1.00 [1.00,1.00] 1.00 [1.00,1.00] 1.00 [1.00,1.00] 1.00 [1.00,1.00] LDECC 0.10 [-0.1,0.30] 0.10 [-0.1,0.30] 0.10 [-0.1,0.30] 0.10 [-0.1,0.30] 0.10 [-0.1,0.30] 0.10 [-0.1,0.30] LDPpa 0.93 [0.85,1.02] 0.90 [0.77,1.00] 1.00 [1.00,1.00] 0.96 [0.91,1.00] 0.93 [0.85,1.00] 1.00 [1.00,1.00] LDP 0.29 [0.29,0.29] 0.17 [0.17,0.17] 1.00 [1.00,1.00] 0.50 [0.50,0.50] 0.33 [0.33,0.33] 1.00 [1.00,1.00] $n=10k$ LD3 1.00 [1.00,1.00] 1.00 [1.00,1.00] 1.00 [1.00,1.00] 1.00 [1.00,1.00] 1.00 [1.00,1.00] 1.00 [1.00,1.00] PC 0.50 [0.17,0.83] 0.50 [0.17,0.83] 0.50 [0.17,0.83] 1.00 [1.00,1.00] 1.00 [1.00,1.00] 1.00 [1.00,1.00] PC 0.00 [0.00,0.00] 0.00 [0.00,0.00] 0.00 [0.00,0.00] 0.47 [0.16,0.78] 0.50 [0.17,0.83] 0.45 [0.14,0.76] LDECC 1.00 [1.00,1.00] 1.00 [1.00,1.00] 1.00 [1.00,1.00] 1.00 [1.00,1.00] 1.00 [1.00,1.00] 1.00 [1.00,1.00] LDECC 0.10 [-0.1,0.30] 0.10 [-0.1,0.30] 0.10 [-0.1,0.30] 0.30 [0.00,0.60] 0.30 [0.00,0.60] 0.30 [0.00,0.60] LDPpa 1.00 [1.00,1.00] 1.00 [1.00,1.00] 1.00 [1.00,1.00] 1.00 [1.00,1.00] 1.00 [1.00,1.00] 1.00 [1.00,1.00] LDP 0.29 [0.29,0.29] 0.17 [0.17,0.17] 1.00 [1.00,1.00] 0.50 [0.50,0.50] 0.33 [0.33,0.33] 1.00 [1.00,1.00] $n=20k$ LD3 0.97 [0.90,1.00] 0.95 [0.85,1.00] 1.00 [1.00,1.00] 1.00 [1.00,1.00] 1.00 [1.00,1.00] 1.00 [1.00,1.00] PC 0.90 [0.70,1.00] 0.90 [0.70,1.00] 0.90 [0.70,1.00] 1.00 [1.00,1.00] 1.00 [1.00,1.00] 1.00 [1.00,1.00] PC 0.00 [0.00,0.00] 0.00 [0.00,0.00] 0.00 [0.00,0.00] 0.10 [-0.1,0.30] 0.10 [-0.1,0.30] 0.10 [-0.1,0.30] LDECC 1.00 [1.00,1.00] 1.00 [1.00,1.00] 1.00 [1.00,1.00] 1.00 [1.00,1.00] 1.00 [1.00,1.00] 1.00 [1.00,1.00] LDECC 0.00 [0.00,0.00] 0.00 [0.00,0.00] 0.00 [0.00,0.00] 0.00 [0.00,0.00] 0.00 [0.00,0.00] 0.00 [0.00,0.00] LDPpa 0.97 [0.90,1.00] 0.95 [0.85,1.00] 1.00 [1.00,1.00] 1.00 [1.00,1.00] 1.00 [1.00,1.00] 1.00 [1.00,1.00] LDP 0.29 [0.29,0.29] 0.17 [0.17,0.17] 1.00 [1.00,1.00] 0.50 [0.50,0.50] 0.33 [0.33,0.33] 1.00 [1.00,1.00]
Table D.8: Performance on the Sachs protein dataset (Sachs et al. 2005) from bnlearn (Scutari 2010). Discrete data were evaluated with $\chi^{2}$ independence tests ($\alpha=0.001$; 95% confidence intervals in brackets).
Sachs Protein Expression SDC = 1: ERK $\to$ AKT SDC = 0: JNK ↛$\not\to$ P38 Time Tests Time Tests Oracle LD3 0.013 37 0.013 37 PC 0.077 740 0.079 740 LDECC 0.079 653 0.081 646 LDP 0.017 46 0.017 52 $n=5k$ LD3 0.005 [0.005,0.006] 37.0 [37.0,37.0] 0.005 [0.005,0.005] 37.0 [37.0,37.0] PC 0.738 [0.729,0.747] 715.4 [708.08,722.72] 0.757 [0.741,0.773] 715.4 [708.08,722.72] LDECC 0.699 [0.688,0.709] 640.6 [638.15,643.05] 0.657 [0.596,0.718] 604.7 [550.92,658.48] LDP 0.006 [0.006,0.006] 47.2 [45.63,48.77] 0.009 [0.004,0.015] 52.4 [51.88,52.92] $n=10k$ LD3 0.007 [0.007,0.007] 37.0 [37.0,37.0] 0.007 [0.006,0.007] 37.0 [37.0,37.0] PC 1.388 [1.374,1.402] 730.4 [726.86,733.94] 1.394 [1.384,1.405] 730.4 [726.86,733.94] LDECC 1.199 [1.045,1.353] 605.1 [528.14,682.06] 1.041 [0.83,1.253] 536.7 [428.01,645.39] LDP 0.008 [0.008,0.008] 46.0 [46.0,46.0] 0.008 [0.008,0.009] 52.0 [52.0,52.0] $n=20k$ LD3 0.011 [0.011,0.011] 37.0 [37.0,37.0] 0.011 [0.01,0.011] 37.0 [37.0,37.0] PC 2.718 [2.689,2.746] 733.6 [732.19,735.01] 2.72 [2.676,2.763] 733.6 [732.19,735.01] LDECC 2.5 [2.451,2.549] 647.4 [644.05,650.75] 2.472 [2.438,2.506] 644.4 [641.0,647.8] LDP 0.012 [0.012,0.013] 46.6 [45.42,47.78] 0.013 [0.013,0.013] 52.0 [52.0,52.0]
Table D.9: Total number of independence tests and runtimes (95% confidence intervals in brackets).
Appendix E Real-World Causal Fairness Experiments
Computing Resources
All causal fairness experiments used an Apple MacBook (M2 Pro Chip; 12 CPU cores; 16G memory).
E.1 COMPAS Recidivism Prediction
Data Preprocessing
We used the data file compas-scores-two-years.csv from ProPublica, [^14] https://github.com/propublica/compas-analysis retaining the following columns for our analyses:
  1. 1.
    race (float64)
  2. 2.
    sex (float64)
  3. 3.
    age_cat (float64)
  4. 4.
    juv_fel_count (int64)
  5. 5.
    juv_misd_count (int64)
  6. 6.
    juv_other_count (int64)
  7. 7.
    priors_count (int64)
  8. 8.
    c_charge_degree (float64)
  9. 9.
    two_year_recid (int64)
  10. 10.
    decile_score (int64)
  11. 11.
    race_binary (int64)
Race was filtered to retain African American ($n=3696$) and white individuals ($n=2454$). For each exposure-outcome pair, $|\mathbf{Z}|=7$. When exposure was race and outcome was decile_score, $\mathbf{Z}$ contained all features other than two_year_recid (as this temporally proceeds decile_score, violating A1). When exposure was race and outcome was two_year_recid, $\mathbf{Z}$ contained all features other than decile_score.
White ($n$ = 2454) African American ($n$ = 3696) Age 37.73 (12.76) 32.74 (10.86) Charge degree (felony) 0.6 (0.49) 0.69 (0.46) Charge degree (misdemeanor) 0.4 (0.49) 0.31 (0.46) General recidivism decile score 3.74 (2.6) 5.37 (2.83) Juvenile felony count 0.03 (0.3) 0.1 (0.51) Juvenile misdemeanor count 0.04 (0.28) 0.14 (0.61) Juvenile other count 0.09 (0.45) 0.14 (0.56) Priors count 2.59 (3.8) 4.44 (5.58) Sex (female) 0.23 (0.42) 0.18 (0.38) Sex (male) 0.77 (0.42) 0.82 (0.38) Two-year recidivism (no) 0.61 (0.49) 0.49 (0.5) Two-year recidivism (yes) 0.39 (0.49) 0.51 (0.5)
Table E.1: Summary statistics for African American versus white individuals in the COMPAS dataset after preprocessing. We report mean values (standard deviations).
Refer to caption Age Refer to caption Sex Refer to caption Charge Degree Refer to caption Priors Count Refer to caption Two-year Recidivism Refer to caption General Recidivism Decile Score
Figure E.1: Distributions for African American versus white individuals in the COMPAS dataset after preprocessing.
Figure E.2: Average number of CI tests performed and average time (seconds) per method on COMPAS experiments.
$DS$$CD$$R$$A$$JF$$PC$ (A) $\alpha=0.005$ $SDC=1$ $WCDE=0.694\;[0.548,0.839]$ $p=0.000$
$DS$$CD$$R$$A$$JF$$PC$ (B) $\alpha=0.01$ $SDC=1$ $WCDE=0.695\;[0.55,0.84]$ $p=0.000$
$DS$$CD$$R$$A$$JF$$PC$$S$$JM$ (C) $\alpha=0.05$ $SDC=1$ $WCDE=0.657\;[0.51,0.804]$ $p=0.000$
$TY$$CD$$A$$PC$$S$ (D) $\alpha=0.005$ $SDC=0$ $WCDE=0.003\;[-0.028,0.033]$ $p=0.87$
$TY$$CD$$A$$PC$$S$$R$$JM$ (E) $\alpha=0.01$ $SDC=1$ $WCDE=0.007\;[-0.024,0.037]$ $p=0.675$
$TY$$CD$$A$$PC$$S$$R$$JF$$JM$ (F) $\alpha=0.05$ $SDC=1$ $WCDE=0.005\;[-0.026,0.035]$ $p=0.766$
Figure E.3: LD3 results on the COMPAS dataset at different significance levels ($\alpha=[0.005,0.01,0.05]$). DAGs represent the predicted parent set for the outcome only, with all other relations abstracted away. (A–C) Exposure race (R; red) and outcome general recidivism decile score (DS; blue). Known parents of outcome DS are in yellow. (D–F) Exposure race and outcome two-year recidivism (TY; blue). A = age; CD = charge degree; DS = decile score; JF = juvenile felony count; JM = juvenile misdemeanor count; PC = priors count; R = race; S = sex; TY = two-year recidivism. Weighted CDE ($WCDE$) is reported with $p$-value ($p$) and 95% confidence intervals in brackets.
$DS$$A$$PC$$R$ (A) $\alpha=0.005$ $SDC=1$ $WCDE=0.709\;[0.564,0.853]$ $p=0.000$
$DS$$A$$PC$$R$ (B) $\alpha=0.01$ $SDC=1$ $WCDE=0.71\;[0.565,0.855]$ $p=0.000$
$DS$$A$$PC$$R$ (C) $\alpha=0.05$ $SDC=1$ $WCDE=0.71\;[0.565,0.855]$ $p=0.000$
$TY$$PC$$R$ (D) $\alpha=0.005$ $SDC=1$ $WCDE=0.058\;[0.029,0.088]$ $p=0.000$
$TY$$PC$$R$ (E) $\alpha=0.01$ $SDC=1$ $WCDE=0.059\;[0.029,0.089]$ $p=0.000$
$TY$$PC$$R$$A$$S$ (F) $\alpha=0.05$ $SDC=1$ $WCDE=0.004\;[-0.026,0.034]$ $p=0.797$
Figure E.4: PC results on the COMPAS dataset at different significance levels ($\alpha=[0.005,0.01,0.05]$). DAGs represent the predicted parent set for the outcome only, with all other relations abstracted away. (A–C) Exposure race (R; red) and outcome general recidivism decile score (DS; blue). Known parents of outcome DS are in yellow. (D–F) Exposure race and outcome two-year recidivism (TY; blue). A = age; DS = decile score; PC = priors count; R = race; S = sex; TY = two-year recidivism. Weighted CDE ($WCDE$) is reported with $p$-value ($p$) and 95% confidence intervals in brackets.
$DS$$A$$PC$$R$ (A) $\alpha=0.005$ $SDC=1$ $WCDE=0.711\;[0.566,0.856]$ $p=0.000$
$DS$$A$$PC$$R$ (B) $\alpha=0.01$ $SDC=1$ $WCDE=0.709\;[0.564,0.853]$ $p=0.000$
$DS$$CD$$R$ (C) $\alpha=0.05$ $SDC=1$ $WCDE=1.559\;[1.396,1.723]$ $p=0.000$
$TY$$A$$PC$ (D) $\alpha=0.005$ $SDC=0$ $WCDE=0.012\;[-0.018,0.042]$ $p=0.445$
$TY$$PC$$R$$A$$S$ (E) $\alpha=0.01$ $SDC=1$ $WCDE=0.006\;[-0.024,0.036]$ $p=0.693$
$TY$$A$$S$$R$ (F) $\alpha=0.05$ $SDC=1$ $WCDE=0.082\;[0.052,0.113]$ $p=0.000$
Figure E.5: LDECC results on the COMPAS dataset at different significance levels ($\alpha=[0.005,0.01,0.05]$). DAGs represent the predicted parent set for the outcome only, with all other relations abstracted away. (A–C) Exposure race (R; red) and outcome general recidivism decile score (DS; blue). Known parents of outcome DS are in yellow. (D–F) Exposure race and outcome two-year recidivism (TY; blue). A = age; CD = charge degree; DS = decile score; PC = priors count; R = race; S = sex; TY = two-year recidivism. Weighted CDE ($WCDE$) is reported with $p$-value ($p$) and 95% confidence intervals in brackets.
E.2 Liver Transplant Allocation
Additional Background
The US liver transplant system has seen several key policy changes designed to optimize organ distribution and improve patient outcomes. In 2002, the model for end-stage liver disease (MELD) was adopted to determine patients’ priority for liver transplants based on the severity of their illness (Malinchoc et al. 2000). In June 2013, UNOS implemented the Share 35 policy to increase access to organs for patients with MELD scores $\geq 35$, given the high waitlist mortality in this subgroup (Wald et al. 2013). In January 2016, the MELD-sodium (i.e., MELD-Na) score incorporated serum sodium levels into the calculation (UNOS 2015). In February 2020, the acuity-circle (AC) policy for distributing livers replaced the longstanding donation service area-based policy (UNOS 2020). The AC policy uses a series of concentric circles, centered on the donor hospital, to determine the distribution range for available livers. In July 2023, the calculation formula for MELD was updated to include additional variables, including female sex and serum albumin (UNOS 2022). The new model, known as MELD 3.0, was expected to further refine the sorting of transplant candidates based on their medical urgency for transplant (Kim et al. 2021).
Data Preprocessing and Feature Selection
We used the National Standard Transplant Analysis and Research (STAR) dataset for patient-level information on waitlisted transplant candidates, recipients, and donors from the Organ Procurement & Transplantation Network (OPTN) up to October 2022 (OPTN 2024). We filtered to include adult patients awaiting or having undergone a whole liver transplant from deceased donors in 2017-2019 and 2020-2022, keeping features relevant for discrimination discovery (Appendix E.2). Time frames were selected as stable periods without changes in the UNOS allocation policy.
We used the following features in the case study. Except for the liver allocation outcome, all features were determined at time of registration or listing. Thus, A1 holds. These features include all the possible parents to the liver allocation outcome available in the OPTN dataset, thus satisfying A2 to the best of our knowledge. For time period 2017-2019, total observations was $n=21,101$ patients. For time period 2020-2022, total observations was $n=22,807$ patients. Below, we provide the definition of each variable included in our causal analysis. Summary statistics are reported in Table E.6.
  1. 1.
    Sex (exposure, protected attribute): Recipient sex.
  2. 2.
    Liver allocation (outcome): Did the candidate receive a liver transplant?
  3. 3.
    Recipient blood type: Recipient blood group at registration.
  4. 4.
    Initial age: Age in years at time of listing.
  5. 5.
    Ethnicity: Recipient ethnicity category.
  6. 6.
    Hispanic/Latino: Is the recipient Hispanic/Latino?
  7. 7.
    Education: Recipient highest educational level at registration.
  8. 8.
    Initial MELD: Initial waiting list MELD/PELD lab score.
  9. 9.
    Active exception case: Was this an active exception case?
  10. 10.
    Exception type: Type of exception relative to hepatocellular carcinoma (HCC).
  11. 11.
    Diagnosis: Primary diagnosis at time of listing.
  12. 12.
    Initial status: Initial waiting list status code.
  13. 13.
    Number of previous transplants: Number of prior transplants that the recipient received.
  14. 14.
    Weight: Recipient weight (kg) at registration.
  15. 15.
    Height: Recipient height at registration.
  16. 16.
    BMI: Recipient body mass index (BMI) at listing.
  17. 17.
    Payment method: Recipient primary projected payment type at registration.
  18. 18.
    Region: Waitlist UNOS/OPTN region where recipient was listed or transplanted.
Unos Policy (2017–2019) Unos Policy (2020–2022) LD3 ($\alpha=0.01$) LD3 ($\alpha=0.05$) LD3 ($\alpha=0.01$) LD3 ($\alpha=0.05$) $\mathbf{A}_{\text{{DE}}}$ active exception case, diagnosis, education, ethnicity, exception type, initial age, initial MELD, region, weight active exception case, diagnosis, education, ethnicity, exception type, initial age, initial MELD, recipient blood type, region, weight active exception case, diagnosis, education, ethnicity, initial age, initial MELD, payment method, region, weight active exception case, diagnosis, education, ethnicity, initial age, initial MELD, payment method, region, weight SDC 1 1 1 1 WCDE 0.033 [0.016,0.049] 0.025 [0.008,0.042] 0.032 [0.018,0.047] 0.032 [0.018,0.047] $p$-value 0.000 0.003 0.000 0.000
Table E.2: Results obtained with nonparametric $\chi^{2}$ independence tests and double machine learning. Total features is 18 (Appendix E.2). For time period 2017-2019, total observations is $n=21101$ patients. For time period 2020-2022, total observations is $n=22807$ patients. With no covariate control, SDC = 1 and the weighted CDE is 0.049 [0.033,0.066] ($p$-value = 0.000).
Unos Policy (2017–2019) Unos Policy (2020–2022) PC ($\alpha=0.01$) PC ($\alpha=0.05$) PC ($\alpha=0.01$) PC ($\alpha=0.05$) $\mathbf{A}_{\text{{DE}}}$ education, recipient blood type None active exception case, intial MELD, recipient blood type active exception case, initial age, recipient blood type, region SDC 1 1 1 1 WCDE 0.046 [0.03,0.063] 0.049 [0.033,0.066] 0.037 [0.023,0.050] 0.028 [0.015,0.041] $p$-value 0.000 0.000 0.000 0.000
Table E.3: SDC assessment and weighted CDE estimation with double machine learning using results from PC Algorithm (Spirtes, Glymour, and Scheines 2001), as implemented by Kalisch and Bühlman (2007). We define $\mathbf{A}_{\text{{DE}}}$ using the union method (PC; Appendix D.2), though results for $\mathbf{A}_{\text{{DE}}}$ were exactly the same using PC.
Unos Policy (2017–2019) Unos Policy (2020–2022) LDECC ($\alpha=0.01$) LDECC ($\alpha=0.05$) LDECC ($\alpha=0.01$) LDECC ($\alpha=0.05$) $\mathbf{A}_{\text{{DE}}}$ payment method payment method, region active exception case, diagnosis, initial age, initial MELD, recipient blood type diagnosis, recipient blood type, region SDC 1 1 1 1 WCDE 0.049 [0.032,0.065] 0.048 [0.032,0.064] 0.032 [0.019,0.046] 0.032 [0.017,0.048] $p$-value 0.000 0.000 0.000 0.000
Table E.4: SDC assessment and weighted CDE estimation with double machine learning using results from LDECC using the union method for selection $\mathbf{A}_{\text{{DE}}}$ (LDECC; Appendix D.2) (Gupta, Childers, and Lipton 2023). Results for $\mathbf{A}_{\text{{DE}}}$ were almost concordant using LDECC, with the exception of 2020-2022 when $\alpha=0.05$ (where the intersection method excluded diagnosis).
LD3 LDECC PC Ratio PC/LD3 Ratio LDECC/LD3 Policy $\alpha$ Time Tests Time Tests Time Tests Time Tests Time Tests 2017-2019 0.01 0.100 75 20.304 3303 237.324 35943 2373.24 479.24 203.04 44.04 0.05 0.101 75 19.947 3188 231.844 36512 2295.48 486.83 197.50 42.51 2020-2022 0.01 0.107 75 47.402 6959 465.278 63225 4348.39 843.00 443.01 92.79 0.05 0.090 71 519.707 69926 528.372 72517 5870.80 1021.37 5774.52 984.87
Table E.5: Total number of independence tests performed and runtimes in seconds for LD3 vs PC (Kalisch and Bühlman 2007) and LDECC (Gupta, Childers, and Lipton 2023) on the liver transplant allocation dataset. While PC is worst-case exponential in time complexity (Spirtes, Glymour, and Scheines 2001), LDECC is worst-case exponential for some structures and polynomial for others (Gupta, Childers, and Lipton 2023). This could explain the wide variation in LDECC runtimes across experimental settings.
Unos Policy (2017-2019) Female (n = 7679) Male (n = 13422) p-value Test Active exception case 0.36 (0.73) 0.48 (0.83) 7.241e-28 t-test Diagnosis 1 (PSC: Primary Sclerosing Cholangitis) 0.03 (0.18) 0.04 (0.2) 0.037 $\chi^{2}$ Diagnosis 6 (AHF: acute hepatic failure) 0.06 (0.23) 0.02 (0.15) 0.004 $\chi^{2}$ Diagnosis 7 (Cancer) 0.09 (0.28) 0.16 (0.37) 0.010 $\chi^{2}$ Height 161.9 (7.46) 175.91 (8.51) 0.000 t-test Initial MELD 20.5 (10.21) 18.83 (9.87) 1.588e-31 t-test Payment method 0.53 (0.5) 0.54 (0.5) 0.012 $\chi^{2}$ Recipient age 54.46 (12.42) 56.03 (10.74) 4.091e-22 t-test Weight 75.97 (18.53) 90.63 (19.59) 0.000 t-test Unos Policy (2020-2022) Female (n = 8574) Male (n = 14233) p-value Test Active exception case 0.39 (0.58) 0.43 (0.63) 3.629e-07 t-test Ethnicity 9 (Multiracial, non-hispanic) 0.01 (0.08) 0.0 (0.07) 0.022 Fisher’s exact Exception type 1 (Unknown) 0.29 (0.45) 0.28 (0.45) 0.022 $\chi^{2}$ Height 161.96 (7.8) 176.36 (8.25) 0.000 t-test Initial MELD 22.13 (10.52) 20.99 (10.48) 1.862e-15 t-test Initial status 0.04 (0.2) 0.02 (0.12) 2.858e-31 t-test Recipient age 53.83 (12.75) 54.74 (11.64) 3.059e-08 t-test Weight 75.76 (18.71) 91.1 (20.32) 0.000 t-test
Table E.6: Mean values (standard deviations) for features with statistically significant differences between males and females ($\alpha=0.05$). Summary statistics for all features are available on GitHub (https://anonymous.4open.science/r/LD3-4440).